### v1

# A Systematic Survey on Large Language Models for Algorithm Design

Abstract.
Algorithm Design (AD) is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and superior solutions. Over the past three years, the integration of LLMs into AD (LLM4AD) has progressed significantly, finding applications in diverse areas such as optimization, machine learning, mathematical reasoning, and scientific exploration. Given the rapid development and broadening scope of this field, a systematic review is both timely and essential. This paper provides a systematic review of the works on LLM4AD. First, we present an overview and summary of existing studies. Then, we present a systematic categorization, and a review of existing works along four dimensions including the role of LLMs, search techniques, prompt strategies, and application fields. We also discuss the achievements and challenges in each area and the capabilities of LLM4AD in addressing them. Finally, we explore current limitations and propose several open questions and promising directions for future research.
Algorithm design, Large language model, Learn to optimize, Optimization, Heuristic, Hyperheuristic, Evolutionary computation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXcopyright: none
1. Introduction
Algorithms play a crucial role in addressing a myriad of problems across diverse fields, including industry, economics, healthcare, and technology (Kleinberg, 2006; Cormen et al., 2022). The traditional hand-crafted approach to algorithm design requires extensive expert knowledge and considerable effort, which is tedious and time-consuming. There has been a growing interest in the integration of learning and computational intelligence techniques to streamline and enhance the algorithm development process (Bengio et al., 2021; Tang and Yao, 2024).
In recent years, Large Language Models (LLMs) have emerged as a significant breakthrough in generative artificial intelligence. LLMs are distinguished by their huge model size, extensive training data, and remarkable performance in various research areas, including mathematical reasoning (Ahn et al., 2024), code generation (Jiang et al., 2024), and scientific discovery (Wang et al., 2023a).
In the past three years, the integration of large language models for algorithm design (LLM4AD) has become an emerging area of research that promises to revolutionize how algorithms are conceived, optimized, and implemented. The impressive power and adaptability of LLMs have shown promising in improving and transforming the algorithm design process, such as heuristic generation (Liu et al., 2024d), code optimization (Ma et al., 2024a), and even the creation of novel algorithms (Girotra et al., 2023) tailored to specific problems. This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the solutions produced (Liu et al., 2024d; Romera-Paredes et al., 2024).
While LLM4AD is gaining traction, systematic reviews in this nascent field remain sparse. Existing literature primarily focuses on isolated applications of LLMs in specific algorithmic contexts, lacking a systematic overview of the methodologies, applications, challenges, and future directions. Several papers have been carried out to survey the LLM for/with specific optimization topic (Huang et al., 2024b; Wu et al., 2024a; He et al., [n. d.]) and others review the general LLM applications (Hadi et al., 2024) or the application on specific problems including electronic design automation (Zhong et al., 2023), planning (Pallagani et al., 2024), soft engineering (Hou et al., 2023), recommendation (Wu et al., 2024b) and agent (Wang et al., 2024c). This paper aims to fill this gap by providing a systematic review of the current state of large language models in algorithm design. We will explore various applications, discuss key challenges, and propose directions for future research. By synthesizing these insights, this paper contributes to a deeper understanding of the potential of LLMs in enhancing algorithm design and lays the groundwork for further innovations in this exciting field. The contributions of this paper are as follows:
  • Systematic Review of LLM4AD: We present a systematic review of the developments in using large language models for algorithm design, covering a significant corpus of 180+ research papers published in the last three years. We not only synthesize the current state of research but also categorize the research works, providing a critical analysis of methodologies, results, and algorithm design advancements. We expect it to be a helpful resource for both new researchers entering the field and experienced experts looking for a consolidated update on recent developments.
  • Development of a Multi-dimensional Taxonomy: We introduce a multi-dimensional taxonomy that categorizes the works and functionalities of LLM4AD into four distinct dimensions: 1) Roles of LLMs in algorithm design, which delineates how these models contribute to or enhance algorithm design, including LLMs as optimizers, LLMs as predictors, LLMs as extractors, and LLMs as designers; 2) Search methods, which explores the various approaches used by LLMs to navigate and optimize search spaces in algorithm design; 3) Prompt methods, which examines how diverse prompting strategies are used; and 4) Application domains, which identifies the key fields and industries where LLMs are being applied to solve complex algorithmic challenges. This taxonomy not only clarifies the landscape but also aids in identifying gaps and opportunities for future research.
  • Discussion on Limitations and Future Directions: We go beyond mere summarization of existing literature to critically analyze the limitations present in current research on LLMs for algorithm design. We discuss challenges such as scalability, interpretability, high cost, and security. Furthermore, we propose potential future research directions that could address these limitations, including developing domain-specific LLMs, exploring multi-modal LLMs, implementing interactive systems with human experts, using LLMs for algorithm assessment and understanding LLM behavior, advancing fully automatic algorithm design, and establishing benchmarking standards for systematic evaluation of LLMs in algorithm design. This discussion is intended to spur novel approaches and foster further advancements in the field.
2. Methodology and Taxonomy
2.1. Scope of Survey
This paper aims to conduct a systematic survey and classification of existing research works in the emerging field of Large Language Model for Algorithm Design (LLM4AD). We further delineate the scope of our survey as follows:
  • We do not intend to cover all literature on LLMs and algorithms. Specifically, we exclude works focused on other branches, such as algorithms for optimizing LLMs, including prompt engineering algorithms (Sahoo et al., 2024) and algorithms for training LLMs (Ackerman et al., 2024).
  • The term Large Language Models refers to language models of sufficient scale to enable robust zero-shot performance across various tasks, including linguistic comprehension, code generation, and mathematical reasoning. These models typically utilize a transformer architecture and operate in an autoregressive manner (Zhao et al., 2023).
    • Studies employing smaller models for algorithm design, such as those prevalent in conventional model-based algorithms and machine learning-assisted algorithms (Bengio et al., 2021), are excluded. While it is challenging to precisely define “large” models, most leading-edge LLMs contain over one billion parameters (Zhao et al., 2023; Minaee et al., 2024).
    • Research utilizing other large models that lack language processing capabilities, such as purely vision-based models, are not considered. However, multi-modal LLMs that include language processing are within our scope.
  • The term Algorithm in this context refers to a set of mathematical instructions or rules designed to solve a problem, particularly when executed by a computer (Cormen et al., 2022). This broad definition encompasses traditional mathematical algorithms (Ahn et al., 2024), most heuristic approaches (Mart et al., 2018; Michel Gendreau, 2019), and certain agents or policies that can be interpreted as algorithms (Xi et al., 2023).
2.2. Statistics
We introduce the detailed pipeline for paper collection and scanning, which consists of four stages:
[…]
Four stages for paper collection.
Figure 1. Four stages for paper collection.
(a) Number of publications
(b) Country distribution
(c) Institutions
[…]
Figure 2. Overview on LLM4AD papers.
  • Stage I Data extraction and collection: We collect the related papers through Google Scholar, Web of Science, and Scopus. The logic of our search is the title must include any combinations of at least one of the following two groups of words “LLM”, “LLMs”, “Large Language Model”, “Large Language Models” and “Algorithm”, “Heuristic”, “Search”, “Optimization”, “Optimizer”, “Design”, “Function” (e.g., LLM and optimization, LLMs and algorithm). As a fast-growing research field, the majority of papers on LLM4AD are published online as pre-print manuscripts such as in Arxiv, which results in a notably more paper collected from Google Scholar (800+) than Web of Science and Scopus. After removing duplicated papers, we ended up with 850 papers as of July 1, 2024.
  • Stage II Abstract Scanning: We first check the title and abstract of each paper to efficiently exclude papers that are not relative. The criteria used for exclusion include these papers that are not in English, not for algorithm design and not using large language models. After scanning, 260 papers are remaining.
  • Stage III Full Scanning: In this stage, we carefully check the entire manuscript of each paper. We remove these papers that do not include relevant content in their main paper. After scanning, there are 160 papers left.
  • Stage IV Supplementation: In order to avoid missing any important contributions, we searched the related works manually according to our past knowledge in this field. After integrating the additional papers, we got 180+ papers in the end.
[…]
Figure 3. The word cloud is generated from the titles and abstracts of all reviewed papers, with each word appearing at least five times. It features the top 80 keywords, organized into four color-coded clusters. The Blue Cluster, which prominently features the keyword “Language”, includes other frequently used terms such as “strategy”, “reasoning”, “prompt”, and “function”, which are crucial for task resolution. The Red Cluster is centered around “GPT”. This cluster also contains terms like “fine-tuning”, “text”, and “accuracy”, which are crucial in model training and inference. The Green Cluster focuses on search and optimization, encompassing terms such as “evolutionary algorithm”, “combination”, and “diversity”. Finally, the Yellow Cluster emphasizes the role of LLMs in scientific discovery, highlighting keywords such as “scientist”, “hypothesis”, and “concept”.
We will first present an overview of the LLM4AD paper list and then present a taxonomy to systematically review the progress. In addition to the organized list of papers, this survey also incorporates some publications released after July 1, 2024. We will update the statistics as needed.
2.3. Overview
Fig. 2(a) illustrates the trend in the number of papers published over time, with the timeline expressed in months. The graph shows a marked rise in research activity related to LLM4AD, particularly noting that most of the studies have been conducted in the last year. This suggests that LLM4AD is an emerging field, and we expect a significant increase in research output in the near future as scholars from diverse fields become aware of its considerable potential. It is worth noting that as a rapidly expanding field, the majority of works are initially published as preprints on Arxiv. Many of these have subsequently been accepted by top conferences such as ICML and NeurIPS.
Fig. 2(c) and Fig. 2(b) display the leading institutions and their respective countries contributing to publications on LLM4AD. The United States leads, closely followed by China, with these two countries alone accounting for 50% of the publications. The next eight countries, including Singapore, Canada, and Japan, collectively contribute one-third of the total publications. Prominent institutions involved in this research include esteemed universities such as Tsinghua University, Nanyang Technological University, and the University of Toronto, alongside major corporations like Huawei, Microsoft, and Google. This distribution underscores the widespread interest in the research topics and their substantial relevance to practical applications in the real world.
In Fig. 3, the word cloud is generated from the titles and abstracts of all reviewed papers, with each word appearing at least five times. It showcases the top 80 keywords, organized into four color-coded clusters on “Language”, “GPT”, “search and optimization”, and “scientific discovery”. Several keywords such as “evolution”, “strategy”, “optimizer”, and “agent” are also highlighted.
[…]
Figure 4. A four-dimensional taxonomy.
2.4. Taxonomy
Research in the field of LLM4AD primarily explores three fundamental components: 1) LLMs, 2) algorithms, and 3) applications. Studies within LLM4AD are categorized based on the variations and interactions of these elements. This paper presents a taxonomy organized into four dimensions: 1) LLM Roles, 2) Search Methods, 3) Prompt Methods, and 4) Applications.
  • LLM Roles: This dimension examines how LLMs are integrated into algorithm development, dividing the workds into four categories: LLMs as Optimizers (LLMaO), LLMs as Predictors (LLMaP), LLMs as Extractors (LLMaE), and LLMs as Designers (LLMaD).
  • Search Methods: This category involves the search methods used. It includes a variety of search strategies, from the basic and widely utilized random search methods to more complex approaches like single-point-based, population-based, and uncertainty-guided methods.
  • Prompt Methods: The effectiveness of pre-trained LLMs is heavily influenced by prompt engineering. We investigate the typical prompt strategies used in LLM4AD including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection prompts.
  • Applications: The broad application of LLM4AD covers diverse fields. We identify the main domains including optimization, machine learning, industry, and scientific discovery.
3. LLM Roles
3.1. LLMs as Optimizers (LLMaO)
In LLMaO, LLMs are utilized as a black-box optimizer within an optimization algorithm to generate and refine solutions (Fig. 5). This innovative use of LLMs extends beyond their traditional applications in natural language processing to more complex problem-solving tasks that involve searching and optimization. The process typically involves employing the LLM in an iterative search framework, where the model actively participates in proposing potential solutions based on feedback received at each step. The integration of LLMs into optimization tasks leverages their ability to understand and generate complex patterns and solutions, which might be infeasible with traditional optimization methods alone.
[…]
Figure 5. Large Language Models as Optimizers (LLMaO). LLMs serve as optimizers within the algorithm to generate new solutions. This typically involves using the LLM in an iterative search process to enhance solution quality. Here, the algorithm and its parameters are usually crafted by humans.
One of the initial efforts to utilize LLMs as optimizers in algorithm design is by Yang et al. (2023). They leverage the in-context learning capabilities of LLMs to generate novel solutions for specific problems, based on previously evaluated solutions. The process of generating solutions is considered a form of in-context learning, where evaluated solutions serve as contextual information. Typically, these solutions are organized to help the LLM understand preferences for optimized outcomes, thereby enabling it to propose superior alternatives. This method is applied iteratively to refine solutions further. Yang et al. (2023) have successfully demonstrated this technique across various domains, including continuous and combinatorial optimization, as well as machine learning tasks.
From an Evolutionary Computation (EC) perspective, using LLMs to generate solutions from existing data can be seen as analogous to search operators in EC algorithms. For instance, Liu et al. (2023a) introduce the use of LLMs as evolutionary operators to tackle multi-objective problems. This method involves breaking down a multi-objective problem into simpler single-objective tasks, with LLMs acting as black-box search operators for each sub-problem, thereby suggesting new solutions. This decomposition framework helps prevent LLMs from being overwhelmed by the complexity of direct multi-objective optimization. In a related study, Liu et al. (2024b) explore the integration of LLMs within evolutionary algorithms (EAs), not just for generating solutions but also for guiding selection, crossover, and mutation processes. Meanwhile, Brahmachary et al. (2024) propose a novel population-based evolutionary framework that includes both exploration and exploitation pools, with solutions being exchanged during the optimization process and LLMs generating solutions for both pools. Differing from direct solution generation, Lange et al. (2024) investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their “EvoLLM” approach, which showed superior performance over baseline algorithms in synthetic BBOB functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms could further improve the performance of EvoLLM. LLMaO has also been used for hyperparameter optimization by Custode et al. (2024), who present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations.
Beyond traditional optimization tasks, LLMs are increasingly being used in prompt engineering to enhance zero-shot performance, a process often referred to as “automatic prompt optimization” (Zhou et al., 2022). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models. Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity, and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors, as explored in studies by Wang et al. (2023b) and Guo et al. (2023). Reflection-based methods, in particular, mimic the trial-and-error process of skilled human prompt designers. Despite the progress, challenges remain, as Ma et al. (2024d) note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. Moreover, even semantically correct reflections do not always result in effective prompts in a single refinement step, often due to unpredictable model responses. To address these issues, they proposed a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models. A recent development in this field is the gradient-based prompt optimization method, termed GPO, introduced by Tang et al. (2024c). This method identifies two critical components of gradient-based search: the update direction and the update method. At each step, GPO retrieves relevant prompts from the optimization trajectory to determine the update direction, then applies a generation-based refinement strategy for the update, while managing the edit distance through a cosine-based decay strategy.
3.2. LLMs as Predictors (LLMaP)
LLMaP employs LLMs as surrogate models, which serve as stand-ins for more complex or computationally expensive models. As illustrated in Fig. 6, these LLMs are utilized to predict the outcomes or responses, functioning either in a classification or regression context (Hao et al., 2024). In classification tasks, the LLMaP approach involves the LLM predicting categorical labels or classes based on input data. On the other hand, in regression tasks, the LLM predicts continuous outputs rather than categories. Compared to other model-based predictors, such as the Gaussian process and conventional neural networks, 1) LLMs are capable of processing and generating human-like responses based on the training they have received on vast datasets. This capability allows them to understand and interpret complex patterns in data, making them suitable for tasks where traditional modeling techniques might struggle due to the intricacies or subtleties in the data (Egami et al., 2024). 2) Pre-trained LLMs can significantly reduce the computational load and time required compared to training high-fidelity models (Jawahar et al., 2023; Hao et al., 2024).
[…]
Figure 6. Large Language Models as Predictors (LLMaP). LLMs are utilized iteratively in algorithms to predict a solution’s outcomes or responses, typically functioning in classification or regression tasks.
The majority of LLMaP works use LLMs as regression models to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by Jawahar et al. (2023). It offers a cost-effective alternative for performance estimation in neural architecture search. Zhang et al. (2024e) introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, Li et al. (2023a) introduce CodonBERT, LLMs for mRNAs, which is used for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and was trained on over 10 million mRNA sequences from various organisms. The model can be used for predicting mRNA properties and outperforms previous methods, as demonstrated on a flu vaccine dataset. Another application on battery electrolytes design by Soares et al. (2024) demonstrates the use of transformer-based models in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (Alipour-Vaezi and Tsui, 2024) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (Zhang et al., 2024a)
Several studies have explored the application of LLMs in classification tasks. For instance, Hao et al. (2024) introduce the LAEA model, which employs LLMs as surrogate models within evolutionary algorithms. This approach is used for both classification, determining the quality of solutions as either good or bad, and regression, by approximating the value of solutions. This method significantly aids in solving optimization problems by evaluating new solutions based on historical data, thus eliminating the need for costly model training. In another study, Chen et al. (2023b) develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance.
Moving beyond binary classification, Bhambri et al. (2024) utilize LLMs to predict discrete actions for constructing reward shaping functions in reinforcement learning (RL). Their method demonstrated effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. Wang et al. (2024d) explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks. Lastly, Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product searches. By finetuning LLMs, they achieved significant improvements, underscoring the adaptability and effectiveness of LLMs in complex search tasks.
Other works study LLM-based classification models. For example, Hao et al. (2024) introduce the LAEA, which utilizes LLMs as surrogate models in evolutionary algorithms and investigates both classification (predicting whether a solution is good or bad), and regression (approximating its value). It helps solve optimization problems by evaluating the quality of new solutions based on historical data eliminating expensive model training. Chen et al. (2023b) discuss a label-free node classification method using LLMs to annotate nodes, which are then used to train graph neural networks for enhanced performance. Except for the binary classification task, Bhambri et al. (2024) use LLMs to predict discrete actions for constructing reward shaping functions in RL, demonstrating effectiveness in the BabyAI environment. Wang et al. (2024d) explore the use of LLMs in federated search, employing them in a zero-shot setting to select resources effectively. Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product search, achieving significant improvements by finetuning LLMs.
LLMs can respond to more than numerical data, which are quite different from conventional prediction models. For example, Egami et al. (2024) utilize LLMs to generate imperfect annotation surrogates for documents in computational social science research, combined with high-quality labels to provide valid statistical inference for downstream analyses. In addition, the multi-modal LLMs demonstrate exceptional capability in utilizing multi-model information. An evolutionary strategy with a vision-language model has been developed by Wong et al. (2024b) for practical 3D car designs. The vision LLM is used to predict practicality degree for designs that satisfy the given text prompt.
3.3. LLMs as Extractors (LLMaE)
LLMaE employs LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then strategically utilized in the enhancement of algorithm-based problem solving (Fig 7). This process leverages the advanced natural language understanding capabilities of LLMs, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional analytical methods (Wu et al., 2024c). By integrating these insights into algorithm design, developers can create more precise and efficient solutions tailored to specific needs.
[…]
Figure 7. Large Language Models as Extractors (LLMaE). LLMs are employed to extract features or specific knowledge from target problem and/or algorithms to enhance problem-solving.
LLMs have demonstrated exceptional capabilities in feature extraction across various domains for algorithm design, surpassing traditional heuristics and conventional deep learning models. For instance, Kristiadi et al. (2024) explore the use of LLMs in accelerating principled Bayesian optimization in the molecular space. LLMs are used as fixed feature extractors for standard Bayesian optimization surrogate models, and parameter-efficient finetuning methods are leveraged to obtain the posterior of the LLM surrogate. The experiments show that LLMs can be useful for Bayesian optimization over molecules, but only if they have been trained or finetuned with domain-specific data.
Beyond feature extraction, LLMs excel in text comprehension and knowledge extraction, when compared to conventional network models. An example in LLM-assisted algorithm selection by Wu et al. (2024c) utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. Extensive experiments validate the performance superiority of the proposed model, showcasing the efficacy of each key module. In another innovative application, Du et al. (2024c) propose a novel approach using a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem and decreasing energy consumption and AI model implementation costs. Additionally, Becker et al. (2024) discuss the use of LLMs in the automotive and supplier industries, specifically focusing on retrieval-augmented generation systems to improve information retrieval from technical documentation. Memduhoğlu et al. (2024) use LLM embeddings to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
The multi-modal capabilities of LLMs also distinguish them from traditional feature extraction methods. Park et al. (2024) demonstrate that incorporating lexical information extracted from LLMs into an acoustic-based speaker diarization system through a probabilistic multi-modal decoding process and joint acoustic and lexical beam searches can significantly improve speech-processing performance.
3.4. LLMs as Designers (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 8). This utilization of LLMs extends their application beyond traditional boundaries, enabling them to actively participate in algorithm development by generating heuristics, writing code snippets, or formulating functions that can be integrated into larger systems. By doing so, LLMs can significantly accelerate the algorithm design process, reduce human effort, and bring creativity and optimization to algorithm development, which is difficult to achieve through conventional methods alone.
[…]
Figure 8. Large Language Models as Designers (LLMaD). LLMs are used to directly create algorithms or specific components, which are commonly incorporated iteratively to continuously search for better designs.
Function design might be one of the most straightforward application domains. Eureka (Ma et al., 2024b) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for reinforcement learning. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. In tests across 29 different reinforcement learning environments, including 10 types of robots, Eureka surpassed human experts in 83% of the tasks, with an average improvement of 52%. Another study on automatic reward function design (Li et al., 2024b) leverages LLMs to automatically design dense reward functions for reinforcement learning agents in environments with sparse rewards, termed Auto MC-Reward. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. FunSearch (Romera-Paredes et al., 2024) adopts LLMs for function generation in an evolutionary framework with a multi-iland population management. FunSearch demonstrated promising results on both mathematical problems and optimization problems.
AEL (Liu et al., 2023b) presents an early attempt to adopt LLM as a designer to design an algorithm idea and its code implementation. It adopts LLM in an evolutionary framework to create, combine, and revise the algorithmic heuristic for the traveling salesman problem. The results show a significant improvement over simple greedy algorithms and robustness across different sizes. EoH (Liu et al., 2024d) is the new version of AEL with five prompt strategies for heuristic space searching. EoH is applied to the key heuristic component design for guided local search. The results outperform human-crafted metaheuristics and deep-learning-based neural solvers. Similar frameworks have been applied to cost-aware acquisition function design for Bayesian optimization (Yao et al., 2024a), adversary attack algorithm design for image adversary attack (Guo et al., 2024c), and heuristic design for edge server task scheduling Yatong et al. (2024). LLaMEA van Stein and Bäck (2024) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (Ye et al., 2024a) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. With the reflection, better results on black-box cases (no problem-specific information is provided in the prompts) have been observed. Moreover, Yan et al. (2024) explore the use of LLMs to generate drafts of high-quality algorithms for electronic health records phenotyping, aiming to improve efficiency and scalability in observational health research and Hazra et al. (2024) combine LLMs with heuristic planning to generate cost-effective plans, showing superior performance in LLM planning approaches.
Unlike previous studies on LLMaD that focus on optimizing a single performance criterion, MEoH (Yao et al., 2024b) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space, which can hardly addressed by conventional MOEAs.
LLM-based agent design has also gained much attention. For example, ADAS (Hu et al., 2024) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents, leveraging the Turing Completeness of programming languages to explore novel prompts, tool use, and control flows. They present a novel algorithm, Meta Agent Search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains, highlighting their robustness and potential for general application. Further studies on LLM-based agent systems are discussed in (Sun et al., 2024a) and (Liu et al., 2024e).
There have been significant advancements in using LLMs for optimization modeling. Ahmed and Choudhury (2024) evaluate the performance of various LLMs in converting linguistic descriptions into mathematical optimization problems. The study introduces LM4OPT, a fine-tuning framework for smaller LLMs that reduces the performance gap between smaller and larger models, with GPT-4 achieving a high accuracy on the NL4Opt dataset without relying on additional named entity information. Similarly, Tang et al. (2024a) identify four critical requirements for the training dataset of operational research LLMs and develops a semi-automated process called OR-Instruct to create synthetic data tailored to these requirements. They introduce the IndustryOR benchmark, the first industrial benchmark for testing LLMs on real-world OR problems. By applying the data from OR-Instruct to various open-source LLMs, the authors demonstrate significantly improved capabilities for optimization modeling. Moreover, Mostajabdaveh et al. (2024) introduce a novel multi-agent modeling framework for translating natural language descriptions into optimization models. Instead of relying on the solver execution, it adopts relations identifier agents and a multi-agent verification mechanism for effective assessment.
4. Search Methods
In this section, we investigate two branches of the interaction of LLMs with search methods: 1) search methods are designed algorithms and 2) search methods are LLM-based searching frameworks for algorithm design. In the first category, search methods are the target of LLMs, examples including LLMs for evolutionary algorithms and Bayesian optimization. In the other category, LLMs are used in the search methods to be iteratively queried to design better algorithms. This section categorizes the works according to the type of search methods without distinguishing the two branches.
4.1. Sampling
The most straightforward search manner is by repeatedly instructing LLM to sample new designs (Zhou et al., 2023). The best sample is selected as the final design. However, recent studies have shown that simple sampling from LLMs is insufficient in algorithm design (Zhang et al., 2024b).
4.1.1. Beam Search
Beam search explores a limited set of promising paths, or “beams”, in the search space rather than only one. Beam search has been widely used in diverse domains. For example, LLM-based prompt engineering (Zhou et al., 2023; Pryzant et al., 2023; Ma et al., 2024d) leads to promising results, e.g., (Zhou et al., 2023) demonstrates cases where when the beam search size increases from 4 to 128, the performance beats the human expert.
Except for prompt engineering, Lin et al. (2023) introduce Text2Motion, a language-based planning framework that enables robots to execute complex sequential manipulation tasks. It adopts beam search over tokens and enhances task planning with LLMs using feasibility heuristics. In contrast, Hazra et al. (2024) adopt beam search over actions. It introduces ”SayCanPay” to integrate the world knowledge of LLMs with heuristic planning principles to produce grounded and efficient action sequences, demonstrating superior performance in extensive evaluations compared to other LLM planning methods. Moreover, Park et al. (2024) use beam search with lexical information from LLMs for enhancing speech-processing tasks and Wei et al. (2024) propose a confidence-aware sub-structure Beam Search as a decoding method operating at the sub-structure level for structured data generation.
4.1.2. MCTS
MCTS is a tree search algorithm that uses random sampling to build a search tree and make decisions based on the results of these samples. It is particularly effective in problems with large and complex search spaces. For example, Dainese et al. (2024) propose GIF-MCTS, a code generation strategy with Generate, Improve and Fix using MCTS, which outperforms baselines on the Code World Models Benchmark and other benchmarks. In GIF-MCTS, nodes in the tree are programs and edges are actions. Each action taken from a parent node produces a new complete program, which is split into a “state” part and a “rollout” part, and stored in a child node. The upper confidence bound for trees is used to select which action to take. Similarly, VerMCTS (Brandfonbrener et al., 2024) constructs a search tree with progressive widening to effectively manage large action spaces defined by lines of code. In this search tree, expansion and evaluation are guided by the verifier, serving as a computationally inexpensive upper bound on the value function in the code synthesis MDP. Moreover, Wang et al. (2023b) regard prompt optimization as a strategic planning challenge, utilizing a principled planning algorithm based on MCTS to effectively explore the expert-level prompt space. Drawing inspiration from human trial-and-error learning, the proposed PromptAgent generates precise insights and detailed instructions by analyzing model errors and providing constructive feedback.
4.2. Single-point-based Search
4.2.1. Hillclimb
Hillclimb search iteratively searches for better results, where the new one is generated based on the last one and the better one is survived. A basic version of Hillclimb is investigated in paper by Zhang et al. (2024b) for algorithm design. It achieves competitive results when compared to population-based evolutionary algorithms. The reasoning over the existing status is usually adopted to enhance each search step. For example, Li et al. (2024b) leverage LLMs to automatically design dense reward functions for reinforcement learning agents in environments with sparse rewards, termed Auto MC-Reward, which iteratively refines the reward function based on feedback from the agent’s interactions with the environment and Yang et al. (2023) develop a multi-module framework with innovative feedback mechanisms, demonstrating through both LLM-based and expert evaluations that LLMs can effectively produce scientific hypotheses that are both novel and reflective of reality.
4.2.2. Neighborhood Search
Neighborhood search methods are optimization techniques that explore the solution space by iteratively modifying a current solution to find an improved one in a structured neighborhood. Notably, the LLM-GS framework (Liu et al., 2024f) has introduced a scheduled hill climbing algorithm that leverages two distinct strategies: 1) Programmatic Neighborhood Generation, which involves generating a neighborhood of programs by selecting a node from the Abstract Syntax Tree (AST) of a given program and substituting it with a subtree that is randomly generated in accordance with the production rules and sampling strategies of a Domain-Specific Language (DSL); and 2) Latent Space Representation, where the neighborhood is defined in a latent space by training a variational autoencoder on a corpus of randomly generated DSL programs, thereby creating a more abstract and potentially more informative neighborhood structure. Additionally,  Jin et al. (2023) contribute to this line of research by defining neighborhood algorithms based on a minimum cosine similarity threshold of 60% between the embeddings of a query vector and code snippets, enhancing the search process by ensuring that algorithmic variations are semantically coherent and aligned with the underlying problem structure.
4.2.3. Gradient-based Search
Gradient-based search leverages a (pseudo) gradient direction in the generation of the new design in each iteration. Unlike searching in continuous space with an explicit differentiable objective formulation, calculating the gradient in the text and code space is challenging. As a result, an estimated descent direction is used as a pseudo gradient in the text and code space.
Pryzant et al. (2023) introduce automatic prompt optimization, which improves prompts automatically by adjusting them based on natural language “gradients” derived from training data. This approach has led to a substantial improvement in LLM performance on various NLP tasks and jailbreak detection by up to 31%. Moreover, Nie et al. (2024) explore the use of LLMs as interactive optimizers for solving maximization problems in a text space through natural language and numerical feedback. LLMs exhibit effectiveness when provided with directional feedback, which is an extension of first-order feedback in the natural language domain. Furthermore, Tang et al. (2024c) introduce a novel Gradient-inspired LLM-based Prompt Optimizer (GPO) that draws parallels with gradient-based model optimizers. By incorporating updated direction and method concepts from gradient-based optimization, GPO enhances prompt optimization strategies, leading to significant performance enhancements in experimental assessments. This integration of LLMs with gradient-based search methods showcases the potential for improving prompt optimization and overall LLM performance.
For an accurate gradient, Guo et al. (2024a) perform the gradient-based search in a mapped continuous space and proposes a collaborative optimization strategy that combines a gradient-based optimizer with an LLM for tackling complex non-convex optimization challenges. The LLM is utilized to derive solutions from natural language instructions, while the gradient-based optimizer performs locally optimal updates at each iteration. This collaborative approach leverages the strengths of both optimizers, with the inferred results from LLMs serving as starting points for subsequent stages of gradient optimization.
4.2.4. Reinforcement Learning
Reinforcement learning learns to make decisions by interacting with an environment to maximize cumulative rewards. Recent works have explored innovative applications and enhancements of this method. For instance, GPTswarm (Zhuge et al., 2024) and the work by Duan et al. (2023) leverage LLMs and reinforcement learning techniques to optimize code, reducing complexity and enhancing efficiency. Similarly, Liu et al. (2024f) present a novel LLM-guided search framework (LLM-GS) aimed at addressing the sample inefficiency in state-of-the-art programmatic reinforcement learning methods by utilizing the programming expertise of LLMs to boost search efficiency.
Additionally, LLMs have been utilized in the design of reinforcement learning systems, particularly in the creation of reward-shaping functions. Bhambri et al. (2024) tackle the issue of sample inefficiency by proposing a framework called MEDIC, which uses LLMs to generate a guide policy for constructing reward-shaping functions. This framework shows substantial improvements in sample complexity for RL agents across various domains in the BabyAI environment.
4.3. Population-based Search
Evolutionary search is the main tool investigated due to its effectiveness and robustness in complex search space (O’Reilly and Hemberg, 2024; Wu et al., 2024a). According to the number of objectives, it can be roughly categorized into single-objective evolutionary search and multi-objective evolutionary search.
4.3.1. Single-objective Evolutionary Search
Evolutionary algorithms are the major solutions when integrating LLM for algorithm design (Liu et al., 2024d; Romera-Paredes et al., 2024). Through maintaining a population of diverse individuals, they usually search the space effectively.
The study by Yang et al. (2023) represents an initial exploration of utilizing LLMs as optimizers for algorithm design. It suggests leveraging the in-context learning capability of LLMs to generate new solutions for target problems based on existing evaluated solutions. Although it does not explicitly mention evolutionary search, the search paradigm behind it can be regarded as a population-based evolutionary search. Liu et al. (2023a) introduce LLMs as evolutionary operators for addressing multi-objective problems by decomposing them into single-objective sub-problems and employing LLMs as black-box search operators for each sub-problem. This decomposition-based approach enables LLMs to tackle multi-objective problems indirectly, overcoming their inherent challenges in comprehending such problems. Additionally, Liu et al. (2024b) explore the use of LLMs as evolutionary operators in conventional EA, guiding them in solution selection, crossover, and mutation processes. Moreover, Brahmachary et al. (2024) propose a population-based evolutionary framework comprising exploration and exploitation pools, with LLMs facilitating solution generation for both pools and enabling communication between them during optimization. Lange et al. (2024) investigate the application of LLMs in designing evolution strategies, introducing a prompting strategy to enhance mean statistic performance.
In contrast to the integration of LLM to generate hard-coded solutions, the effective evolutionary search of prompt and code generation poses challenges due to the lack of suitable diversity measurement, selection, and population management methods on the complex search space. The majority of workers adopt a greedy way (Liu et al., 2024d; Chen et al., 2024a; van Stein and Bäck, 2024; Li et al., 2024b; Narin, 2024; Guo et al., 2024c; Yao et al., 2024a). In these works, a population of individuals is maintained and only the ones with better fitness will survive. Romera-Paredes et al. (2024) adopt a multi-island evolutionary algorithm to explore a diverse population with a dynamically adjusted population size. Moreover, Wong et al. (2024a) explore the potential of leveraging generative artificial intelligence models to create digital artifacts for creative and engineering applications, utilizing an multi-factorial evolutionary algorithm to drive an LLM. The proposed LLM2FEA approach demonstrates the ability to generate novel and practical designs, as evidenced by experimental results in the context of 3D aerodynamic design.
To enable effective diversity control over the population, Lehman et al. (2023) and Hu et al. (2024) have investigated quality diversity methods (Pugh et al., 2016). Specifically, they utilize the MAP-Elites algorithm with a grid of niches to generate diverse and high-quality solutions. A pre-existing solution is evaluated and placed into the map, with subsequent iterations perturbing solutions within inhabited niches. New candidate solutions are evaluated and assigned niches based on behavior characterization, becoming champions if they outperform current inhabitants or fill unfilled niches. This process gradually fills the map with improved solutions over time.
4.3.2. Multi-objective Evolutionary Search
Liu et al. (2023a) propose MOEA/D-LMO to use LLM to solve continuous multi-objective optimization problems. Benefiting from the decomposition-based framework, the in-context learning of LLM is used to generate new solutions for each subproblem. For multi-objective LLM instruction generation, Yang and Li (2023) develop InstOptima, which utilizes an LLM to simulate instruction operators and improve instruction quality to simultaneously optimize performance, length, and perplexity of the instruction. Moreover, Morris et al. (2024) introduce guided evolution for the development of neural architectures. It adopts NSGA-II and optimizes both accuracy and model size. MEoH (Yao et al., 2024b) introduce a dominance-dissimilarity score for multi-objective heuristic design. In contrast to conventional MOEAs, it achieves a good balance of exploration and exploitation in the heuristic search space.
4.4. Uncertainty-guided Search
Uncertainty-guided search harnesses the advantages of BOED and BO frameworks within the search process incorporating LLMs. It first establishes a prior that describes the initial belief about the parameters of interest and then utilizes uncertainty-driven strategies over the derived posterior to iteratively update the belief, ultimately achieving efficient decision-making during the search process in a sample-efficient manner.
This search paradigm with uncertainty has demonstrated its superiority in some scenarios. For example, in tasks of eliciting human preferences information using LLMs (Li et al., 2023b; Piriyakulkij et al., 2023), Handa et al. (2024) introduce OPEN that takes advantage of LLMs and Bayesian Optimal Experiment Design (BOED). Specifically, LLMs are utilized to extract environment-relevant features and translate abstract feature queries into natural language questions to interact with users through conversations, BOED is employed to identify the most informative query for efficiently learning preference information. Experimental studies have demonstrated the superiority of the OPEN approach in content recommendation tasks when compared to both LLM- and BOED-based preference elicitation methods. To improve the multi-turn and decision-theoretic reasoning capability over a set of arbitrary items, Austin et al. (2024) formulate the natural language-based preference elicitation within a Bayesian optimization framework. The proposed PEBOL algorithm maintains a probabilistic prior over user preferences in the form of Beta distributions. The generation of queries is facilitated by LLM-based acquisition functions incorporating strategies like Thompson Sampling and Upper Confidence Bound, which aim to learn user preferences efficiently balancing exploration and exploitation. In the domain of LLM decoding, Grosse et al. (2024) propose a probabilistic model named ULTS. This method frames the decoding process as a tree-search problem and identifies an optimal path in the LLM search tree by sequential decision-making under uncertainty. It firstly establishes a prior belief on the softmax outputs of the LLM and then utilizes the implied posterior samples to guide the search process in a non-myopic fashion.
Among these search methods mentioned in this section, evolutionary algorithms (EAs) are the most used search methods in LLM-based algorithm design (Cai et al., 2024), Table 1 summarizes and compares different EA-based methods in terms of their algorithmic components, i.e., prompt strategy, search strategy, and population management. For prompt strategy, we identify One-Shot prompting (OS), Few-Shot prompting (FS) (Brown et al., 2020), Chain-of-Thought (CoT) (Wei et al., 2022), and Reflection (RF) (Brown et al., 2020). For search methods, we identify Genetic Algorithm (GA), Differential Evolution (DE), Evolutionary Strategy (ES), Multi-objective Evolutionary Algorithm (MOEA) and others. The results reveal that: 1) FS and CoT are the most commonly used prompt engineering strategies. 2) The majority of works use a simple genetic algorithm with greedy population management. 3) LLMaO and LLMaD are more commonly investigated than the other two categories.
Table 1. Existing LLM-based Methods with Evolutionary Computation for Algorithm Design
Approaches Method Prompt Strategy Search Strategy Population Management
LLMaP LAEA (Hao et al., 2024) FS, CoT GA Greedy
LLMaO EvoLLM (Lange et al., 2024) FS ES All
EvolPrompt (Guo et al., 2023) FS, CoT GA, DE Greedy, Fixed Size
EvolPrompting (Chen et al., 2024a) FS, CoT GA Greedy
Instoptima (Yang and Li, 2023) FS, CoT MOEA Greedy
MOEA/D-LMO (Liu et al., 2023a) FS MOEA MOEA/D, Fixed Size
LMEA (Liu et al., 2023a) CoT GA Greedy
LEO (Brahmachary et al., 2024) FS, CoT Two-pop GA Greedy
GE (Morris et al., 2024) FS, CoT MOEA Greedy
LLMaD ELM (Lehman et al., 2023) FS MAP-Elites MAP-Elites
AEL (Liu et al., 2023b) FS, CoT GA Greedy
Eureka (Ma et al., 2024b) FS, RF GA (1+1) Greedy
FunSearch (Romera-Paredes et al., 2024) FS Multi-island GA Greedy, Dynamic Size
EoH (Liu et al., 2024d) FS, OS, CoT GA Greedy, Fixed Size
L-AutoDA (Guo et al., 2024c) FS, OS, CoT GA Greedy
EvolCAF (Yao et al., 2024a) FS, OS, CoT GA Greedy
ReEvo (Ye et al., 2024a) CoT, RF GA Greedy, Fixed Size
1+1 (Zhang et al., 2024b) OS GA (1+1) Greedy, Fixed Size
LLaMEA (van Stein and Bäck, 2024) CoT GA Greedy
Auto MC-Reward (Li et al., 2024b) CoT GA (1+1) Greedy
EROM (Narin, 2024) FS, RF GA (1+1) Greedy
5. Prompt Strategies
(a) Pre-trained and domain LLMs
(b) Prompt engineering techniques
(c) LLMs
[…]
Figure 9. LLM types and promote engineering methods
Fig. 9(a) depicts the percentage of domain or pre-trained LLMs used in the literature. Among them, over 80% chose to use the pre-trained model without any specific finetuning, and about 10% fine-tuned the pre-trained model on domain datasets (Cummins et al., 2024) in which 4.4% are trained from scratch. Fig. 9(c) illustrates the top-used LLMs. LLM4AD papers show a strong preference for GPT models. GPT-4 and GPT-3.5 are the most-used LLMs, which in total take around 50%. Llama-2 are the most used open-source LLM. The development of an effective domain LLMs poses challenges to most algorithm designers. We will provide more discussions in the future work section.
Once we have the pre-trained LLMs, prompt engineering is significant for effectively integrating LLMs in algorithm design. We introduce the main prompt engineering methods including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection (Sahoo et al., 2024) used in LLM4AD papers. Fig. 9(b) shows the number of papers involving different prompt engineering techniques. Few-shot, zero-shot, and chain-of-thought are the most commonly used ones. Many works involve more than one prompt engineering technique.
5.1. Zero-Shot
Zero-Shot (ZS) learning enables a model to comprehend and perform tasks without prior specific training on those tasks. In algorithm design, zero-shot prompting allows direct requests to the LLM for solutions or responses, examples including conversational replies (Sekulić et al., 2024), assertion design (Miftah et al., 2024), reinforcement learning guidance (Zhang et al., 2024e), and molecule generation (Ye, 2024). The model leverages its pre-trained knowledge to generate these responses. Although this method is advantageous for swiftly producing solutions based on general knowledge, it may not consistently deliver tailored responses to the nuanced demands of complex algorithmic challenges (Zhang et al., 2024b).
5.2. Few-Shot
Few-Shot (FS) learning involves providing the model with a handful of examples to illustrate the task before it attempts to solve a similar problem. This approach enhances the model’s understanding of the context and its ability to tailor responses to specific tasks. In algorithm design, this method proves particularly effective by presenting the model with examples of algorithms (Liu et al., 2023b), solutions (Yang et al., 2024), prompts (Wang et al., 2023b; Tang et al., 2024c), and codes (Ma et al., 2024a; Morris et al., 2024). The examples used in the prompt may be manually-designed seeds (Romera-Paredes et al., 2024; Hu et al., 2024) or derived from LLM-generated designs (Liu et al., 2024d; Wong et al., 2024a; Ma et al., 2024c; Morris et al., 2024; Wang et al., 2023b; Tang et al., 2024c; Custode et al., 2024). These samples are typically sorted to suggest a preference to enhance performance (Yang et al., 2024; Romera-Paredes et al., 2024).
5.3. Chain-of-Thought
Chain-of-Thought (CoT) prompting encourages the model to articulate intermediate steps or reasoning paths that lead to the final answer. This technique is particularly valuable in algorithm design, where understanding the step-by-step process or engaging in instructed reasoning over existing designs helps prevent outlier results and fosters effective algorithm development. For instance, Liu et al. (2024d) introduce several prompting strategies, one of which directs the LLM to reason over existing heuristics, summarize the general pipeline, and design a new one based on this reasoning. Custode et al. (2024) prompt the LLM to evaluate whether the current step size is sufficient for convergence and to suggest adjustments based on its reasoning. In addition, multiple LLM-based agents are employed within a CoT framework by Sun et al. (2024b) to enhance heuristic design for a boolean satisfiability problem solver.
5.4. Self-consistency
Self-consistency involves generating multiple answers or solutions to the same prompt and then synthesizing them to improve accuracy and reliability. In algorithm design, this could mean prompting the model multiple times for different approaches to solve a problem and then comparing these solutions to identify the most efficient or robust algorithm. This approach leverages the model’s ability to generate diverse solutions and builds a more comprehensive understanding of possible strategies. For example, Guan et al. (2023) set the number of responses as 10 and uses the majority vote to get the predicted label. As each response may provide different keywords or regular expressions, it takes the union of the keywords or regular expressions to create a candidate set. Moreover, Lehman et al. (2024) sample multiple revised codes from the same selected code and updates the population accordingly.
5.5. Reflection
Reflection in prompt engineering involves asking the model to critique or evaluate its own responses or solutions. After generating an algorithm, the model can be prompted to reflect on the efficiency, potential flaws, or improvements. This not only helps in refining the algorithm but also in understanding deeper nuances, such as space-time trade-offs, which are crucial for advanced algorithm design. Ma et al. (2024d) investigate both implicit and explicit reflection in LLM-driven prompt optimization, which lets LLM analyze the errors and generate a reflection or feedback regarding the current prompt. An additional study by Ye et al. (2024b) incorporates short-term and long-term reflections in heuristic design, while Zhang et al. (2024c) focus on self-evaluation and reasoning over results from network detection. Furthermore, both Ma et al. (2024b) and Narin (2024) adopt LLM for RL reward function design with a reward reflection that monitors the scalar values of all reward components and the task fitness function at various policy checkpoints during training.
6. Applications
6.1. Optimization
In this subsection, we focus on discussing the practical applications of LLM in algorithm design for optimization. We categorize the existing literature into combinatorial optimization, continuous optimization, Bayesian optimization, prompt optimization, and optimization modeling based on the specific types of optimization tasks. We then proceed to compare the various roles played by LLMs, the prompt strategies utilized, and the specific problems or tasks to which they are applied. This comparative analysis is summarized in Table 2, where we list the names of the frameworks or methods proposed by the authors. For studies that do not explicitly name their methods, we assign appropriate designations in our article and denote them with asterisks for easy reference (e.g., MH-LLM* for (Sartori et al., 2024)).
6.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), the design of effective algorithm heuristics has been a significant area of interest for a long time. The emergence of LLMs has now provided new avenues for automatically generating CO algorithms for diverse application tasks.
The Traveling Salesman Problem (TSP) stands out as one of the most renowned combinatorial optimization challenges, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Given its NP-hard nature, heuristic algorithms are commonly employed to tackle this problem. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (Liu et al., 2023b), ReEvo (Ye et al., 2024a) and EoH (Liu et al., 2024d). Differently, OPRO (Yang et al., 2024) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (Liu et al., 2024c) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (Huang et al., 2024c) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance.
In addition to routing problems, the Bin Packing Problem (BPP) is another well-studied problem with diverse applications. The objective of the BPP is to pack items of varying sizes into the smallest number of fixed-sized bins. FunSearch (Romera-Paredes et al., 2024) utilizes an evolutionary approach to evolve heuristics for the BPP. The study showcases that FunSearch is capable of uncovering enhanced heuristics when compared to traditional strategies such as first fit and best fit in addressing the challenges posed by the BPP. In comparison to FunSearch, EoH (Liu et al., 2024d) designs heuristics that not only achieve competitive performance but also demonstrate superior generalization across various problem sizes and capacities, requiring significantly fewer queries. Furthermore, in addition to the previously discussed research, numerous important combinatorial optimization problems have been extensively studied to validate the effectiveness of their proposed methods. Examples include the cap set problem in FunSearch (Romera-Paredes et al., 2024), Flow Job Shop Scheduling Problem (FSSP) in EoH (Liu et al., 2024d), and social networks problem in MH-LLM* (Sartori et al., 2024).
6.1.2. Continuous Optimization
Here we delve into works that deal with optimization problems featuring continuous design variables, encompassing both single- and multi-objective optimization. Specific tasks or benchmarks include synthetic functions as well as real-world applications.
In the realm of single-objective optimization, LLaMEA (van Stein and Bäck, 2024) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (Hansen and Ros, 2010) within IOHexperimenter benchmarking tool (de Nobel et al., 2024), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (Lange et al., 2024) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (Yang et al., 2024) illustrates that LLMs effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (Brahmachary et al., 2024) devises an explore-exploit policy using LLMs for solution generation, the method has been tested and validated in both single- and multi-objective benchmark functions, as well as industrial engineering problems including supersonic nozzle shape optimization, heat transfer, and windfarm layout optimization. In a slight departure from the aforementioned methods that directly employ LLMs for solution generation, LAEA (Hao et al., 2024) introduces LLM-based surrogate models for regression and classification tasks to assess the quality of new solutions for selection. This approach has been validated on 2D test functions using nine mainstream LLMs.
Compared to single-objective continuous optimization, multi-objective optimization problems (MOPs) seek to identify a set of optimal solutions, referred to as Pareto optimal solution sets. Multi-Objective Evolutionary Algorithms (MOEAs) serve as effective tools for addressing MOPs across various domains. An initial exploration of utilizing LLMs to tackle MOPs is introduced in MOEA/D-LMO (Liu et al., 2023a). Benefiting from the decomposition-based framework, the in-context learning process of LLMs is easily incorporated to generate candidate solutions for each subproblem derived from the original MOP. In the realm of large-scale MOPs, LLM-MOEA* (Singh and Bali, 2024) showcases the inferential capabilities of LLMs in multi-objective sustainable infrastructure planning problem. The study highlights the LLM’s proficiency in filtering crucial decision variables, automatically analyzing the Pareto front, and providing customized inferences based on varying levels of expertise. Additionally, CMOEA-LLM (Wang et al., 2024b) leverages LLMs with evolutionary search operators to address the challenges of constrained MOPs and exhibits robust competitiveness in DAS-CMOP test suite (Fan et al., 2020).
6.1.3. Bayesian Optimization
Bayesian optimization (BO) is a model-based optimization paradigm for solving expensive optimization problems and has found wide application in various real-world scenarios (Frazier and Wang, 2016). It typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to select potential solutions in a sample-efficient manner. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (Zhang et al., 2023) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. This approach has demonstrated comparably or better performance when compared with random search and Bayesian optimization on HPOBench dataset (Eggensperger et al., 2021). Furthermore, LLAMBO incorporates LLM capabilities to enhance the efficiency of model-based BO, in which three specific enhancements throughout the BO pipeline including zero-shot warmstarting, surrogate models and candidate point sampler, as well as the end-to-end performance have been systematically investigated, the evaluated tasks are selected from Bayesmark (Uber, 2020) and HPOBench (Eggensperger et al., 2021). Instead of utilizing LLM for direct solution generation, BO-LIFT* (Ramos et al., 2023) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (Dinh et al., 2022) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (Yao et al., 2024a) introduces a novel paradigm to design acquisition functions automatically for cost-aware Bayesian optimization, the approach showcases remarkable efficiency and generalization on synthetic functions and hyperparameter tuning task in JAHS-Bench-201 data set (Bansal et al., 2022). FunBO (Aglietti et al., 2024) discoveres novel and well-performing AFs for BO by extending FunSearch (Romera-Paredes et al., 2024), the discovered AFs are evaluated on various global optimization benchmarks in and out of the training distribution and HPO tasks for RBF-based SVM and AdaBoost algorithms.
6.1.4. Prompt Optimization
Prompt optimization aims to identify the most effective task prompt that maximizes the performance of the LLM on a specific task dataset. Previous studies have primarily focused on discrete optimization for natural language prompts (Deng et al., 2022; Zhang et al., 2022; Xu et al., 2022; Prasad et al., 2022) or continuous optimization in the embedding space (Li and Liang, 2021; Lester et al., 2021). Nevertheless, despite of requiring specialized training for each specific task, these approaches typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (Zhou et al., 2022) utilizes LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (Honovich et al., 2022) and Big-Bench Hard (BBH) (Suzgun et al., 2022). OPRO (Yang et al., 2024) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on GSM8K (Cobbe et al., 2021) and BBH. Inspired by the numerical gradient descent method, APO (Pryzant et al., 2023) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. The outcomes showcase substantial enhancements with no hyperparameter tuning or model training on four NLP benchmark classification tasks, namely Jailbreak, Ethos (Mollas et al., 2020), Liar (Wang, 2017) and Sarcasm (Farha and Magdy, 2020). Similarly, Tang et al. (2024c) introduce a gradient-inspired LLM-based prompt optimizer named GPO and evaluate its performance across complex reasoning, knowledge-intensive, and common NLP tasks on datasets including BBH, GSM8K, MMLU (Hendrycks et al., 2020), WSC (Levesque et al., 2012) and WebNLG (Gardent et al., 2017). Guo et al. (2024b) present a collaborative optimization approach that integrates a gradient-based optimizer and an LLM-based optimizer in an interleaved fashion, the combined method achieves consistent improvements compared with baseline prompt tuning methods on NLU and image classification tasks. By harnessing the benefits of evolutionary algorithms that that strike a good balance between exploration and exploitation, Guo et al. (2023) introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts. The superiority of EvoPrompt over manually designed prompts and previous automatic prompt generation approaches has been demonstrated across 31 datasets, encompassing tasks related to language understanding, generation, and the BIG-Bench benchmark.
6.1.5. Optimization Modeling
Optimization modeling (Berry and Houston, 1995) aims to construct a mathematical model of a real-world optimization problem in a standard form represented with an objective function and a set of constraints (AhmadiTeshnizi et al., 2023; Tang et al., 2024a). In addition to designing effective algorithms to solve optimization problems, the process of modeling or formulating the problems also presents a significant challenge. Even when employing the same optimization algorithm, different modeling formulation may lead to significant differences in solving efficiency. Besides, creating an efficient formulation often demands expertise in optimization skills and specialized domain knowledge in a wide range of applications (Wasserkrug et al., 2024; Saghafian et al., 2015; Shakoor et al., 2016).
Recently, LLMs have opened up promising avenues for automating optimization modeling, facilitating the formulation of problems and the implementation of solutions, thereby democratizing expert knowledge. AhmadiTeshnizi et al. (2023) propose OptiMUS, an LLM-based agent designed to formulate and solve optimization problems interpreted from natural language. The performance of the method is evaluated on a proposed novel dataset named NLP4LP (AhmadiTeshnizi et al., 2023) (Natural Language Processing for Linear Programming), which includes 41 linear programming (LP) and 11 mixed integer linear programming (MILP) problems. Despite its outstanding performance, the heavy reliance on sensitive data submission to proprietary LLMs can pose data privacy concerns in industrial applications. To solve this issue, Tang et al. (2024a) propose training open-source LLMs for optimization modeling that are applicable to real-world industries. They develop and implement OR-Instruct, a semi-automated process for generating synthetic data customized to meet specific requirements. The best-performing resulting model, ORLM, demonstrates state-of-the-art performance on NL4Opt (Ramamonjison et al., 2023), MAMO (Huang et al., 2024a), and the proposed first industrial benchmark, IndustryOR. Ahmed and Choudhury (2024) develop a progressive fine-tuning framework named LM4OPT to enhance LLMs’ specificity for optimization problem formulation task on NL4Opt dataset. Wasserkrug et al. (2024) develop Decision Optimization CoPilot (DOCP), which is accessible to a broad range of decision-makers lacking problem-dependent modeling skills. This tool aids in decision-making by interacting with the user in natural language and utilizes human’s knowledge and feedback to create and solve a problem-specific optimization model. The experiments are conducted targeting three presented requirements for DOCP on facility location problems.
6.1.6. Other Applications
In addition to optimization applications that can be categorized into various fields as mentioned above, we have also discovered applications of LLMs in other optimization-related domains, which may intersect to varying degrees with the aforementioned classifications. For ease of reference, we have categorized them as other applications.
Wu et al. (2024c) propose AS-LLM, which harnesses the representation capabilities of LLMs to extract algorithm features for algorithm selection tasks. The approach integrates a feature selection module to identify critical features and leverages the similarity between algorithm and problem representations. In addition to demonstrating robust performance advantages on the ASlib (Algorithm Selection Library) benchmark (Bischl et al., 2016) in experimental studies, the authors also offer a rigorous upper bound on the model complexity of AS-LLM. Ma et al. (2024a) introduce LLaMoCo, the first instruction-tuning framework tailored to adapt general-purpose LLMs for optimizing code generation. Benefiting from a well-formatted instruction set and a novel two-phase training strategy, the fine-tuned LLMs exhibit robust and superior optimization performance, reducing the requirement for extensive domain knowledge and effort in prompt design. StrategyLLM (Gao et al., 2023) employs four LLM-based agents - the strategy generator, executor, optimizer, and evaluator - in a collaborative manner. The approach enables LLMs to reason inductively and deductively, resulting in the formulation of more generalizable and consistent few-shot prompts compared to CoT prompting methods. Experimental results demonstrate the effectiveness and reliability of StrategyLLM across four challenging tasks, encompassing math reasoning, commonsense reasoning, algorithmic reasoning, and symbolic reasoning.
Table 2. An Overview of Optimization Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Prompt Strategy Specific Problems or Tasks
Combinatorial Optimization AEL (Liu et al., 2023b) LLMaD FS, CoT TSP
ReEvo (Ye et al., 2024a) LLMaD CoT, RF TSP
OPRO (Yang et al., 2024) LLMaO FS TSP
LMEA (Liu et al., 2024c) Mixed FS TSP
MLLM (Huang et al., 2024c) Mixed ZS, FS CVRP
FunSearch (Romera-Paredes et al., 2024) LLMaD FS Cap Set Problem, Online BPP
EoH (Liu et al., 2024d) LLMaD FS, OS, CoT TSP, Online BPP, FSSP
MH-LLM* (Sartori et al., 2024) LLMaD ZS, FS Social Networks Problem
Continuous Optimization LLaMEA (van Stein and Bäck, 2024) LLMaD CoT BBOB
EvoLLM (Lange et al., 2024) LLMaO FS BBOB, Neuroevolution
OPRO (Yang et al., 2024) LLMaO FS Linear Regression
LEO (Brahmachary et al., 2024) LLMaO FS Numerical Benchmarks, Industrial Engineering Problems
LAEA (Hao et al., 2024) FS, CoT LLMaP Ellipsoid, Rosenbrock, Ackley, Griewank
MOEA/D-LMO (Liu et al., 2023a) LLMaO FS ZDT, UF
LLM-MOEA* (Singh and Bali, 2024) LLMaE ZS Multi-objective Sustainable Infrastructure Planning Problem
CMOEA-LLM (Wang et al., 2024b) LLMaO FS DAS-CMOP
Bayesian Optimization HPO-LLM* (Zhang et al., 2023) LLMaO FS HPOBench
LLAMBO (Liu et al., 2024a) Mixed FS, ZS, CoT Bayesmark, HPOBench
BO-LIFT* (Ramos et al., 2023) LLMaD FS Catalyst Optimization
EvolCAF (Yao et al., 2024a) LLMaD FS, OS, CoT Synthetic Functions, JAHS-Bench-201
FunBO (Aglietti et al., 2024) LLMaD FS Synthetic Functions, HPO
Prompt Optimization APE (Zhou et al., 2022) LLMaO FS Instruction Induction, BBH
OPRO (Yang et al., 2024) LLMaO FS GSM8K, BBH, MultiArith, AQuA
APO (Pryzant et al., 2023) LLMaO ZS, FS NLP Benchmark Classification Tasks
GPO (Tang et al., 2024c) LLMaO FS BBH, GSM8K, MMLU, WSC, WebNLG
MaaO (Guo et al., 2024b) LLMaO FS NLU Tasks, Image Classification Tasks
EvoPrompt (Guo et al., 2023) LLMaO CoT, FS Language Understanding and Generation Tasks, BBH
Optimization Modeling DOCP (Wasserkrug et al., 2024) LLMaO FS, CoT Facility Location Problems
OptiMUS (AhmadiTeshnizi et al., 2023) Mixed ZS, FS NL4LP
ORLMs (Tang et al., 2024a) LLMaD ZS NL4Opt, MAMO, IndustryOR
LM4OPT (Ahmed and Choudhury, 2024) LLMaD FS, ZS NL4Opt
Other Applications AS-LLM (Wu et al., 2024c) LLMaE ZS Algorithm Selection
LLaMoCo (Ma et al., 2024a) LLMaD FS Optimization Code Generation
StrategyLLM (Gao et al., 2023) Mixed ZS, FS, CoT Math & Commonsense & Algorithmic & Symbolic Reasoning
6.2. Machine Learning
Table 3. An Overview of Machine Learning Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Prompt Strategy Specific Problems or Tasks
Task Planning Zero-Planner (Huang et al., 2022) LLMaP ZS, FS VirtualHome Tasks
SayCan (Ahn et al., 2022) LLMaP FS Real-world Kitchen Tasks with Robots
LLM-GM (Huang et al., 2023) LLMaP FS Real-world Kitchen Tasks with Robots
Text2Motion (Lin et al., 2023) LLMaP FS Long Sequence Tasks for Robots
ProgPrompt (Singh et al., 2023) LLMaD FS, COT VirtualHome Tasks
SayCanPay (Hazra et al., 2024) LLMaP FS VirtualHome Tasks, Ravens Tasks, BabyAI Tasks
Reinforcement Learning LFG (Shah et al., 2023) LLMaP FS, CoT ObjectNav Tasks, Real-world Tasks
SLINVIT (Zhang et al., 2024e) LLMaP ZS, FS ALFWorld Tasks, InterCode Tasks, BlocksWorld Tasks
MEDIC (Bhambri et al., 2024) LLMaP ZS BabyAI Tasks
Eureka (Ma et al., 2024b) LLMaD FS, CoT IsaacGym Tasks, Bidexterous Manipulation Tasks
EROM (Narin, 2024) LLMaD ZS, CoT IsaacGym Tasks
LLM-MOE (Du et al., 2024c) LLMaP FS, CoT Intelligent Networks
Neural Architecture Search EvoPrompting (Chen et al., 2023a) LLMaD FS, ZS, CoT MNIST Dataset, CLRS Algorithmic Reasoning
HS-NAS (Jawahar et al., 2023) LLMaP FS, ZS, CoT Machine Translation Tasks
LLMatic (Nasir et al., 2023) LLMaD FS, ZS, CoT CIFAR-10 Dataset, NAS-bench-201 Benchmarks
LAPT (Zhou et al., 2024a) LLMaD ZS, FS NAS201, Trans101, DARTs
LLM-GE (Morris et al., 2024) LLMaD FS, ZS, CoT CIFAR-10 Dataset
Graph Learning LLM-Critical (Mao et al., 2024) LLMaD FS, ZS, CoT Critical Node Identification
LLM-GNN (Chen et al., 2023b) LLMaP FS, CoT Label-free Node Classification
ReStruct (Chen et al., 2024c) LLMaE FS, ZS Meta-structure Discovery
AutoAlign (Zhang et al., 2024d) LLMaE FS, ZS Entity Type Inference
KSL (Feng et al., 2023b) LLMaP FS Knowledge Search
Dataset Labeling Inter-Classier (Chiquier et al., 2024) LLMaO FS, ZS iNaturalist Datasets, KikiBouba datasets
DataSculpt (Guan et al., 2023) LLMaP FS Label Function Design
Other Applications LLM2FEA (Wong et al., 2024a) LLMaP FS, CoT Objective-oriented Generation
tnGPS (Zeng et al., 2024) LLMaD FS, OS, CoT Tensor Network Structure Search
L-AutoDA (Guo et al., 2024c) LLMaD FS, OS, CoT Adversarial Attack
In this subsection, we investigate the applications of LLMs in the machine learning domain, focusing on their contribution to algorithmic design. Our analysis begins with a focus on significant application areas, notably task planning and reinforcement learning—key benchmarks in the pursuit of generalized artificial intelligence. Subsequently, we probe into specialized applications including neural architecture search, graph-based learning, and the automation of dataset labeling. While it is not feasible to discuss every application exhaustively, we feature additional uses of LLMs in tasks like tensor network structure search and adversarial attack algorithm generation. These applications are summarized in Table 3.
6.2.1. Task Planning with LLMs
The advent of LLMs has garnered significant interest in their application to planning tasks, owing to their remarkable generative capabilities of instructions in the form of natural language. Pioneering work by Huang et al. (2022) demonstrates the effectiveness of LLMs in leveraging world knowledge to produce actionable plans for VirtualHome environments, showcasing their potentiality in such contexts. This method was further enhanced by Ahn et al. (2022), who employed a value function to ground the world knowledge, enabling the derivation of an optimal policy for robotic task execution in lifelike kitchen settings. In subsequent research, Huang et al. (2023) broaden the scope of grounded information in a more adaptive way. Moreover, Lin et al. (2023) utilize greedy search in combination with heuristic algorithms to refine the planning system. Distinctly, Singh et al. (2023) tackle planning challenges by integrating programming languages, thereby generating code to streamline planning processes. A more comprehensive approach, named SayCanPay (Hazra et al., 2024), is developed for planning by considering both the affordance function and reward within the task framework.
6.2.2. Reinforcement Learning.
Reinforcement Learning (RL) has been the de facto standard for sequential decision-making tasks, and recently, the synergy between RL and LLMs has emerged as a novel trend in the domain. This convergence mirrors the dynamics of task planning, yet places RL at the core of its methodology. Many of the LLM4AD papers on RL is for automatically designing the reward functions (Bhambri et al., 2024; Ma et al., 2024b; Narin, 2024). In addition, Shah et al. (2023) investigate the employment of LLMs for heuristic planning to steer the search process within RL frameworks. Zhang et al. (2024e) integrate LLMs into RL by introducing a Kullback-Leibler divergence regularization term that aligns LLM-driven policies with RL-derived policies. LLMs have also extended their reach to multi-agent RL scenarios, as shown by Du et al. (2024c), who illustrates their application within a Mixture-of-Experts system to direct RL models in the realm of intelligent network solutions.
6.2.3. Neural Architecture Search
Neural Architecture Search (NAS) has emerged as a significant focus within the AutoML community, with recent enhancements attributed to the capabilities of LLMs. LLMs are most commonly used for designing neural architectures. For example, Chen et al. (2024a) have integrated LLMs with evolutionary search to successfully generate code for diverse tasks, including NAS for the MNIST dataset. Nasir et al. (2023) introduce a quality-diversity algorithm tailored to NAS, producing architectures for CIFAR-10 and NAS-bench-201 benchmarks. Moreover, Morris et al. (2024) introduce guided evolution for the development of neural architectures and has been the first to suggest the concept of the evolution of thought in algorithm design. LLMs have been also used for predicting the performance of the network. Jawahar et al. (2023) employ LLMs in predicting NAS performance, combining this approach with evolutionary search to effectively create novel network architectures. Different from the LLM-based architecture design and score prediction, (Zhou et al., 2024a) investigate how to adopt LLM for design principle transfer to reduce and direct the search space. A notable search efficiency improvement is achieved when compared to existing transfer methods for NAS.
6.2.4. Graph Learning
Graph learning constitutes a significant field within machine learning. With the advancing capabilities of LLMs in symbolic reasoning and graph processing, their integration into graph learning has emerged as a recent trend. For example, Chen et al. (2023b) apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both Mao et al. (2024) and Chen et al. (2024c) adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. Zhang et al. (2024d) introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and Feng et al. (2023b) develop the knowledge search language, which leverages LLMs to effectively conduct searches within knowledge graphs.
6.2.5. Dataset Labeling
LLMs have been used for mining semantic and multi-modal information from datasets. LLMs are emplyed to train interpretable classifiers to extract attributes from images (Chiquier et al., 2024) and to generate label functions for weakly supervised learning (Guan et al., 2023).
6.2.6. Other Applications
Other Applications of LLMs extend to a myriad of machine learning tasks, though it is impractical to enumerate them all. Notably, Wong et al. (2024a) capitalize on multi-task learning and the iterative refinement of prompts to foster innovative design approaches. Zeng et al. (2024) integrate LLMs with evolutionary search to develop a heuristic function aimed at efficiently sifting through candidate tensor sets representing the primal tensor network. Furthermore, Guo et al. (2024c) have blazed a trail in employing LLMs to generate novel decision-based adversarial attack algorithms, thus opening up a new diagram for the automatic assessment of model robustness.
6.3. Science Discovery
This subsection is dedicated to exploring scientific discoveries related to algorithm designs. Table. 4 summarizes the related works in this domain. For an in-depth discussion and review of LLMs in scientific discovery, readers are referred to survey papers (Birhane et al., 2023; Hadi et al., 2024; Jablonka et al., 2023).
6.3.1. General Science Discovery
In the realm of general scientific discovery, several frameworks have been developed that leverage LLMs. Notably, Du et al. (2024b) introduce LLM4ED, a framework that employs iterative strategies, including a black-box optimizer and evolutionary operators, to generate and optimize equations. This approach has shown significant advancements in stability and usability for uncovering physical laws from nonlinear dynamic systems. Similarly, Shojaee et al. (2024) present LLM-SR, which combines LLMs’ extensive scientific knowledge and code generation capabilities with evolutionary search. This framework excels in proposing and refining initial equation structures based on physical understanding, outperforming traditional symbolic regression methods in discovering physically accurate equations across multiple scientific domains. Another innovative framework, SGA, introduced by Ma et al. (2024c), is a bilevel optimization framework that merges the abstract reasoning capabilities of LLMs with the computational power of simulations. This integration facilitates hypothesis generation and discrete reasoning with simulations for experimental feedback and optimization of continuous variables, leading to improved performance in discovering constitutive laws and designing molecules.
6.3.2. Chemistry, Biology&Physics
In the fields of chemistry and biology, LLMs are predominantly applied in diverse molecule design scenarios. These include general molecule design (Ishida et al., 2024; Bhattacharya et al., 2024), drug molecule design (Ye, 2024), chemical reactions (Ranković and Schwaller, 2023), and catalyst design (Wang et al., 2024a). Physics has been less studied in LLM4AD papers, Pan et al. (2024) demonstrate the capability of LLMs to perform complex analytic calculations in theoretical physics using multi-step prompt templates.
6.3.3. Mechanics
MechAgents (Ni and Buehler, 2024) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, Du et al. (2024a) adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to state-of-the-art models. Another example in fluid dynamics has been investigated by Zhu et al. (2024), which introduces FLUID-LLM, a novel framework that integrates pre-trained LLMs with spatiotemporal-aware encoding to enhance the prediction of unsteady fluid dynamics, addressing the challenge of complex geometries in computational fluid dynamics. Furthermore, Buehler (2023) study LLM-based methods for forward and inverse mechanics problems including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.
Table 4. An Overview of Science Discovery Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Specific Problems or Tasks
General Scientific Equation Discovery Bilevel (Ma et al., 2024c) LLMaD Physical Scientific Discovery
LLM-SR (Shojaee et al., 2024) LLMaD Scientific Equation Discovery
LLM4ED (Du et al., 2024b) LLMaD Equation Discovery
Chemical ChatChemTS (Ishida et al., 2024) LLMaD Molecule design
Debjyoti Bhattacharya, et al. (Bhattacharya et al., 2024) LLMaO Molecule design
Agustinus Kristiadi, et al. (Kristiadi et al., 2024) LLMaE Molecule design
BoChemian (Ranković and Schwaller, 2023) LLMaE Chemical reaction
Multi-modal MoLFormer (Soares et al., 2024) LLMaP Battery electrolytes formulation design
CataLM (Wang et al., 2024a) LLMaP catalyst design
Gavin Ye (Ye, 2024) LLMaD drug design
DrugAssist (Ye et al., 2023) LLMaO drug design
Biology MLDE (Tran and Hy, 2024) LLMaP Protein design
X-LoRA (Buehler and Buehler, 2024) LLMaP Protein design
CodonBERT (Li et al., 2023a) LLMaP mRNA design and optimization
Revisiting-PLMs (Hu et al., 2022) LLMaP protein function prediction
Mechanics FLUID-LLM (Zhu et al., 2024) LLMaP computational fluid dynamics
MechAgents (Ni and Buehler, 2024) LLMaD mechanics design
MeLM (Buehler, 2023) LLMaP, LLMaD carbon nanotubes and proteins design
Mengge Du, et al. (Du et al., 2024a) LLMaD nonlinear dynamics equation discovery
6.4. Industry
This subsection explores the transformative impact of LLMs on algorithm design across various industries. As we envision the future of manufacturing, LLMs are poised to play a pivotal role. Existing developments, such as those from SymphonyAI  [^1] https://www.symphonyai.com/, focus on enhancing research and investment to refine these models while assessing their performance in terms of safety, bias, and utility. LLMs can be seen as a gateway between humans and machines, facilitating informed decision-making.
For instance, LLMs are instrumental in building a comprehensive intelligent 6G network system (Long et al., 2024), addressing challenges such as low latency, high-frequency bands, high bandwidth, high transmission rates, and overall intelligence (Zhou et al., 2024b). Moreover, LLMs have the potential to revolutionize the telecom industry by streamlining tasks and reducing the need for extensive manpower and engineering expertise. However, it is crucial to address existing challenges to fully realize their potential (Maatouk et al., 2023).
Recently, the application of LLMs in electronic design automation (EDA) has emerged as a promising area of exploration. LLM-based solutions have been developed to enhance interactions with EDA tools (Fang et al., 2024). In particular, within the VLSI design flow, it is essential to verify that the implementation conforms to its specifications. To achieve this, SystemVerilog Assertions (SVA) need to be generated automatically. Generative AI techniques, such as LLMs, can be effectively utilized to facilitate this process (Miftah et al., 2024).
To enhance the reliability and availability of cloud services, conducting root cause analysis (RCA) for cloud incidents is essential. RCACopilot (Chen et al., 2024b) effectively matches incoming incidents with the appropriate incident handlers based on their alert types. It aggregates critical runtime diagnostic information, predicts the incident’s root cause category, and provides an explanatory narrative—all without the need for manual investigations.
More applicable scenarios include various industrial design challenges, such as chip design (Chang et al., 2024), car shape design (Wong et al., 2024b), aircraft concept design (Lovaco et al., 2024) UI design (Feng et al., 2023a), and robot design (Yu et al., 2023). LLMs primarily leverage billions of web resources. However, practical applications require fine-tuning with thousands of specific representations of design intent to recognize the distinct patterns associated with each design task.
Other traditional industrial tasks, such as itinerary planning, can also be enhanced by leveraging the power of LLMs (Xie et al., 2024). LLM-based open-domain urban itinerary planning (Tang et al., 2024b) combines spatial optimization with LLMs to create detailed urban itineraries tailored to users’ requests. This approach extends the scope of itinerary planning by generating personalized urban itineraries based on user requests expressed in natural language.
7. Challenges of LLMs in Algorithm Design
The application of LLM for algorithm design is not without limitations. This section explores the boundaries of LLMs concerning their scalability, inference costs, and their ability to generate novel ideas specifically in the context of designing algorithms.
7.1. Scalability
One of the primary limitations of LLMs in algorithm design is their scalability, particularly for handling long inputs and outputs. LLMs are limited by a fixed context window size, restricting the amount of information they can process at once, which is a challenge for complex algorithmic tasks requiring detailed specifications and extensive data. This limitation not only affects their understanding but also the length of the output they can generate, often necessitating multiple iterative queries for comprehensive solutions. Additionally, even when input and output lengths are sufficient, LLMs struggle with comprehension and reasoning over long inputs, impacting their effectiveness in designing complex algorithms. For instance, LLMs tend to solve only low-dimensional problems when used as optimizers (Yang et al., 2023), and they are typically employed to design a function or heuristic component of an algorithm rather than the entire algorithm when used as designers (Liu et al., 2024d).
7.2. Interpretability
Interpretability is another challenge when using LLMs for algorithm design. It is difficult to trace how specific inputs lead to particular outputs in LLMs. This opacity can be problematic, especially in critical applications, e.g., some industry applications, where understanding the decision-making process is essential for trust and reliability. Efforts to enhance the interpretability of LLMs involve techniques such as feature visualization, model simplification (Liu et al., 2023a), and the study of the contribution of different parts of the data to the model’s behavior (Allen-Zhu and Li, 2023). Despite these efforts, achieving high levels of interpretability without compromising the performance of LLMs remains a key area of ongoing research in the field of artificial intelligence.
7.3. Inference Cost
The deployment of LLMs in algorithm design also faces challenges related to costs, both in terms of training and use. Training domain LLMs requires significant computational resources, often involving extensive GPU or TPU usage over several weeks or even months, which incurs high financial and time costs. Once trained, the inference cost—the computational expense involved in applying the model to new data—also becomes a critical factor, especially when the model is used frequently or requires rapid responses, such as the evolutionary search used in many LLM4AD works (Romera-Paredes et al., 2024). For practical applications in algorithm design, where multiple iterations and tests might be necessary, these costs can accumulate, making the use of LLMs potentially less viable for researchers without substantial computational budgets.
7.4. Security
Integrating LLMs into algorithm design introduces significant security risks, including the mishandling of personal data and the potential creation of harmful code. LLMs are trained on extensive datasets that often contain sensitive information, posing a risk of privacy breaches if these models inadvertently generate outputs influenced by this data. Additionally, the autonomous nature of LLMs can lead to the unintentional development of flawed or malicious algorithms, a critical concern in sectors like financial services and personal data management (Zhao et al., 2024). To address these issues, it is vital to establish stringent data governance and model training protocols that prevent the inclusion of sensitive data, rigorously test algorithms under various conditions to prevent unintended behaviors, and continuously update security measures to counteract emerging vulnerabilities.
7.5. Generation of Novel Ideas
Despite the numerous works on LLM4AD, their ability to produce truly novel ideas is still under scrutiny. LLMs excel at interpolation within the range of their training data but may struggle with extrapolation to new, unseen scenarios. In the context of algorithm design, this means that while LLMs can effectively suggest algorithms or solutions that closely resemble or incrementally improve upon existing ones, they may not be as effective at pioneering radically new algorithmic concepts or methods that deviate significantly from the training data. This limitation is crucial for fields that thrive on innovation and where the generation of novel, out-of-the-box ideas is paramount. The reliance on large volumes of training data also raises concerns about the model’s dependency on historical data, which may bias its outputs towards conventional methods and potentially stifle innovation. While LLM does show the capability of generating new ideas in some tasks (Yao et al., 2024a), how to continuously explore practical novel ideas remains a challenge (Hu et al., 2024).
8. Future Directions
8.1. Domain LLM for Algorithm Design
Instead of using a general pre-trained LLM with linguistic and code generation capability, it is worthwhile studying how to train an LLM specifically for automatic algorithm design tasks.
The following aspects can be explored in developing domain LLMs:
  • Training domain LLMs are costly and resource-consuming. On one hand, advanced light fine-tuning techniques, e.g., LORA, could be used for efficient domain adaptation. On the other hand, the size of algorithm LLMs for a specific application can be reduced with the help of domain data and knowledge.
  • Generating and collecting domain data for algorithm design poses challenges. Different from general code generation or linguistic process tasks, there is no large and formatted data specific for algorithm design. Some Github reps on algorithms might be helpful, however, the resources should be scanned and cleaned for algorithm design. Lehman et al. (2024) provide an example of generation domain data for LLMs, where the samples generated during searching are used as the training dataset to fine-tune a pre-trained model resulting in better performance on the target problem.
  • Instead of learning a text and code generation model, how to learn the high-level algorithm design ideas and the algorithmic revision and reasoning ability remain unexplored. The main challenge is how to promote the automatic reasoning capability of LLMs in inference (Zelikman et al., 2022; Wang and Zhou, 2024), which is also recognized as one of the key technologies behind state-of-the-art LLMs, such as OpenAI o1 (OpenAI, 2024).
8.2. Multi-modal LLM for Algorithm Design
The primary focus of the works discussed in this paper is on utilizing the text comprehension and generation capabilities of LLMs, whether in language, code, or statistics. One distinct advantage of LLMs compared to traditional model-based optimization is their ability to process multi-modal information like humans. Some attempts have been conducted to showcase the advantages of incorporating multi-modal information in algorithm design, as evidenced by Huang et al. (2024c) and Narin (2024). It is anticipated that more principled approaches utilizing multi-modal LLMs will be developed, with further application studies expected in the future.
8.3. Interaction with Human Experts
Further research is needed to explore the interaction between LLMs and human experts in algorithm design (Li et al., 2024a). For example, in LLMaD works, LLMs can be seen as intelligent agents, making it feasible for human experts to step in and take over tasks such as generating, modifying, and evaluating algorithms. Investigating how to facilitate efficient and productive collaboration between LLMs and human experts would be valuable. Ideas and techniques in collective intelligence (Malone and Bernstein, 2022) can be used for this purpose.
8.4. LLM-based Algorithm Assessment
LLMs can be helpful in algorithm assessment. Some attempts have been carried out to automatically evaluate algorithms and assess algorithm design. For example, Sterbini and Temperini (2024) show that LLMs can be used to analyze algorithm descriptions and compare them to the task description given. Such text similarity correlates well both with the grades given by the teacher and with the grades given by the students. OMNI-EPIC (Faldor et al., 2024) utilizes foundation models to automatically produce code that defines the next engaging and learnable tasks. By generating environments and reward functions, OMNI-EPIC has the capability to create a wide range of simulated learning tasks for algorithm evaluation. We expect more studies on LLM-based algorithm assessment.
8.5. Understand the Behavior of LLM
In the majority of works, LLM works as a black-box model. The interpretation of LLM’s behavior not only enriches our understanding of LLM’s behavior but also benefits cases where it is challenging or costly to directly request LLMs. Some attempts have been made to approximate and understand the in-context learning behavior of LLM in solution generation. For example, Liu et al. (2023a) has designed a white-box linear operator to approximate the results of LLM in multi-objective evolutionary optimization. In spite of these preliminary attempts, how to interpret LLM’s behavior is still an open question in many algorithm design cases including heuristic generation and idea exploration.
8.6. Fully Automatic Algorithm Design
Fully automatic algorithm design faces two primary challenges: 1) the generation of novel algorithmic ideas and 2) the creation of complex, lengthy code. While the generation of new ideas has been investigated in some works (Liu et al., 2024d), the complete design of algorithms (instead of a heuristic component), including both heuristic components and detailed code implementations, remains a challenge. Existing applications typically focus on automating components within predefined algorithm frameworks rather than creating new algorithms from scratch. Future research needs to tackle these complexities to advance the field of fully automatic algorithm design.
8.7. Benchmarking LLM4AD
Benchmarking allows a fair, standardized, and convenient comparison and identifying best practices enabling generating new insights for innovation. While we are glad to witness the emergence of diverse research works and applications, LLM4AD still lacks a benchmarking on either algorithm design test sets for systematic and scientific assessment or a strict pipeline and evaluation criteria on large language models for algorithm design. Several attempts have been made on related topics such as mathematical reasoning (Lin et al., 2024) and planning (Valmeekam et al., 2024). (Veličković et al., 2022) presents an algorithm test set built on some basic algorithms and (McLeish et al., 2024) demonstrates the superior performance of LLM when compared to neural algorithm reasoners. In the future, more benchmarks are expected and they are going to play a crucial role in advancing LLM4AD.
9. Conclusion
This paper has provided a systematic and up-to-date survey on large language models for algorithm design (LLM4AD), an area that stands at the confluence of artificial intelligence and computational problem-solving. By systematically reviewing a significant corpus of recent literature, this paper not only highlights the current state and evolution of LLM applications in algorithm design but also introduces a novel multi-view taxonomy with comprehensive lists of papers that categorize the roles and functionalities of LLMs in this domain. This taxonomy serves as a framework for both academic and industrial researchers to understand and navigate the complex landscape of algorithm design using LLMs.
The discussion presented in this paper sheds light on the limitations and challenges currently faced by the field, such as issues of scalability, interpretability, and security. Moreover, the proposed future research directions aim to inspire and guide subsequent studies including the need for domain-specific models, multimodal approaches, interactive systems, and comprehensive benchmarking standards.
As we look forward, it is clear that the intersection of LLMs and algorithm design holds promising potential for revolutionizing how algorithms are developed and implemented. The ability of LLMs to automate and innovate within the algorithm design process could lead to more efficient, effective, and creative algorithms for complex problems across various domains. We hope this paper can contribute to a deeper understanding of this potential and set the stage for future innovations and collaborations in this promising avenue of research.