### v2

# A Systematic Survey on Large Language Models for Algorithm Design

Abstract.
Algorithm Design (AD) is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. Over the past three years, the integration of LLMs into AD (LLM4AD) has progressed significantly, finding applications in diverse areas such as optimization, machine learning, mathematical reasoning, and scientific discovery. Given the rapid development and broadening scope of this field, a systematic review is both timely and essential. This paper provides a systematic review of the works on LLM4AD. First, we present an overview and summary of existing studies. Then, we present a systematic taxonomy and a review of existing works along four dimensions, including the role of LLMs, search techniques, prompt strategies, and applications, with a discussion of the potential and achievements of using LLMs. Finally, we explore current challenges and propose several open questions and promising directions for future research.
Algorithm design, Large language model, Optimization, Heuristic, Hyperheuristic, Evolutionary algorithm.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXcopyright: none
1. Introduction
Algorithms play a crucial role in addressing various problems across diverse fields, including industry, economics, healthcare, and technology (Kleinberg, 2006; Cormen et al., 2022). The traditional hand-crafted approach to algorithm design requires extensive expert knowledge and considerable effort, which is tedious and time-consuming. There has been a growing interest in the integration of learning and computational intelligence techniques to streamline and enhance the algorithm development process (Bengio et al., 2021; Tang and Yao, 2024).
In recent years, Large Language Models (LLMs) have emerged as a significant breakthrough in generative artificial intelligence. LLMs are distinguished by their huge model size, extensive training data, and remarkable performance in various research areas, including mathematical reasoning (Ahn et al., 2024), code generation (Jiang et al., 2024), and scientific discovery (Wang et al., 2023a).
In the past three years, the integration of Large Language Models for Algorithm Design (LLM4AD) has become an emerging area of research that promises to revolutionize how algorithms are conceived, optimized, and implemented. The impressive power and adaptability of LLMs have shown promising in improving and transforming the algorithm design process, such as heuristic generation (Liu et al., 2024d), code optimization (Ma et al., 2024a), and even the creation of novel algorithms (Girotra et al., 2023) tailored to specific problems. This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the produced solutions (Liu et al., 2024d; Romera-Paredes et al., 2024).
While LLM4AD is gaining traction, systematic reviews in this nascent field remain sparse. Existing literature primarily focuses on applications of LLMs in specific algorithmic contexts. For example, several papers have been carried out to survey the LLM for/with specific optimization topic (Huang et al., 2024b; Wu et al., 2024a; He et al., [n. d.]) and others review the general LLM applications (Hadi et al., 2024) or the application on specific problems including electronic design automation (Zhong et al., 2023), planning (Pallagani et al., 2024), soft engineering (Hou et al., 2023), recommendation (Wu et al., 2024b) and agent (Wang et al., 2024c). This paper aims to fill this gap by providing a systematic review with a multi-dimensional taxonomy of the current state of LLMs in algorithm design. We will explore various applications, discuss key challenges, and propose directions for future research. By synthesizing these insights, this paper contributes to a deeper understanding of the potential of LLMs in enhancing algorithm design and lays the groundwork for further innovations in this exciting field. We expect it to be a helpful resource for both new researchers entering the field and experienced experts looking for a consolidated and systematic update on current developments. The contributions of this paper are as follows:
  • Systematic Review of LLM4AD: We present the first systematic review of the developments in using LLMs for algorithm design, covering a significant corpus of 180+ highly related research papers published in the last three years.
  • Development of a Multi-dimensional Taxonomy: We introduce a multi-dimensional taxonomy that categorizes the works and functionalities of LLM4AD into four distinct dimensions: 1) Roles of LLMs in algorithm design, which delineates how these models contribute to or enhance algorithm design; 2) Search methods, which explores the various approaches used by LLMs to navigate and optimize search spaces in algorithm design; 3) Prompt methods, which examines how diverse prompting strategies are used; and 4) Application domains, which identifies the key fields and industries where LLMs are being applied to solve complex algorithmic challenges. This taxonomy not only clarifies the landscape but also aids in identifying gaps and opportunities for future research.
  • Discussion on Challenges and Future Directions: We go beyond mere summarization of existing literature to critically analyze the limitations present in current research on LLMs for algorithm design. Furthermore, we propose potential future research directions, including developing domain-specific LLMs, exploring multi-modal LLMs, facilitating human-LLM interaction, using LLMs for algorithm assessment and understanding LLM behavior, advancing fully automatic algorithm design, and benchmarking for systematic evaluation of LLMs in algorithm design. This discussion is intended to spur novel approaches and foster further advancements in the field.
2. Methodology and Taxonomy
2.1. Scope of Survey
This paper aims to conduct a systematic survey and classification of existing research works in the emerging field of Large Language Model for Algorithm Design (LLM4AD). We do not intend to cover all the literature on LLMs and algorithms. We delineate the scope of our survey as follows:
  • The term Large Language Models refers to language models of sufficient scale. These models typically utilize a transformer architecture and operate in an autoregressive manner (Zhao et al., 2023). 1) Studies employing smaller models for algorithm design, such as those in conventional model-based and machine learning-assisted algorithms (Bengio et al., 2021), are excluded. While it is challenging to precisely define “large” models, most leading-edge LLMs contain over one billion parameters (Zhao et al., 2023; Minaee et al., 2024). 2) Research utilizing other large models that lack language processing capabilities, such as purely vision-based models, are not considered. However, multi-modal LLMs that include language processing are within our scope.
  • The term Algorithm in this context refers to a set of mathematical instructions or rules designed to solve a problem, particularly when executed by a computer (Cormen et al., 2022). This broad definition encompasses traditional mathematical algorithms (Ahn et al., 2024), most heuristic approaches (Mart et al., 2018; Michel Gendreau, 2019), and certain agents or policies that can be interpreted as algorithms (Xi et al., 2023).
We introduce the detailed pipeline for paper collection and scanning, which consists of four stages:
[…]
Four stages for paper collection.
Figure 1. Four stages for paper collection.
  • Stage I Data Extraction and Collection: We collect the related papers through Google Scholar, Web of Science, and Scopus. The logic of our search is the title must include any combinations of at least one of the following two groups of words “LLM”, “LLMs”, “Large Language Model”, “Large Language Models” and “Algorithm”, “Heuristic”, “Search”, “Optimization”, “Optimizer”, “Design”, “Function” (e.g., LLM and optimization, LLMs and algorithm). After removing duplicated papers, we ended up with 850 papers as of July 1, 2024.
  • Stage II Abstract Scanning: We check the title and abstract of each paper to efficiently exclude irrelevant papers. The criteria used for exclusion include these papers that are not in English, not for algorithm design and not using large language models. After scanning, 260 papers are remaining.
  • Stage III Full Scanning: We thoroughly review each manuscript to exclude papers lacking relevant content. After scanning, there are 160 papers left.
  • Stage IV Supplementation: We append some related works manually according to our past knowledge in this field to avoid missing any important contributions. After integrating the additional papers, we got 180+ papers in the end.
(a) Number of publications
(b) Country distribution
(c) Institutions
[…]
Figure 2. Overview on LLM4AD papers.
[…]
Figure 3. The word cloud is generated from the titles and abstracts of all reviewed papers, with each word appearing at least five times. It features the top 80 keywords, organized into four color-coded clusters.
We will first present an overview of the LLM4AD paper list and then present a taxonomy to systematically review the progress. In addition to the organized list of papers, we also incorporate some important publications released after July 1, 2024.
2.2. Overview
Fig. 2(a) illustrates the trend in the number of papers published over time, with the timeline expressed in months. The graph shows a marked rise in research activity related to LLM4AD, particularly noting that most of the studies have been conducted in the last year. This suggests that LLM4AD is an emerging field, and we expect a significant increase in research output in the near future as scholars from diverse fields become aware of its considerable potential.
Fig. 2(c) and Fig. 2(b) display the leading institutions and their respective countries contributing to publications on LLM4AD. The United States leads, closely followed by China, with these two countries alone accounting for 50% of the publications. The next eight countries, including Singapore, Canada, and Japan, collectively contribute one-third of the total publications. Prominent institutions involved in this research include esteemed universities such as Tsinghua University, Nanyang Technological University, and the University of Toronto, alongside major corporations like Huawei, Microsoft, and Google. This distribution underscores the widespread interest in the research topics and their substantial relevance to practical applications in the real world.
In Fig. 3, the word cloud is generated from the titles and abstracts of all reviewed papers, with each word appearing at least five times. It showcases the top 80 keywords, organized into four color-coded clusters on “language”, “GPT”, “search and optimization”, and “scientific discovery”. Several keywords such as “evolution”, “strategy”, “optimizer”, and “agent” are also highlighted.
2.3. Taxonomy
This paper presents a taxonomy organized into four dimensions: 1) LLM Roles, 2) Search Methods, 3) Prompt Methods, and 4) Applications.
  • LLM Roles: This dimension examines how LLMs are integrated into algorithm development, dividing the workds into four categories: LLMs as Optimizers (LLMaO), LLMs as Predictors (LLMaP), LLMs as Extractors (LLMaE), and LLMs as Designers (LLMaD).
  • Search Methods: This category involves the search methods used including the basic sampling methods and more complex approaches like evolutionary algorithms, and uncertainty-guided methods.
  • Prompt Methods: The effectiveness of pre-trained LLMs is heavily influenced by prompt engineering. We investigate the typical prompt strategies used in LLM4AD including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection prompts.
  • Applications: The broad application of LLM4AD covers diverse fields. We identify the main domains including optimization, machine learning, industry, and scientific discovery.
3. LLM Roles
3.1. LLMs as Optimizers (LLMaO)
In LLMaO, LLMs are utilized as a black-box optimizer within an algorithm framework to generate and refine solutions (Fig. 4). The integration of LLMs into optimization tasks leverages their ability to understand and generate complex patterns and solutions and good flexibility in prompt engineering (Yang et al., 2024; Liu et al., 2023a, 2024c), which might be infeasible for traditional optimizers. However, due to their black-box nature, they usually lack interpretability and face challenges in large-scale generation (Liu et al., 2023a).
[…]
Figure 4. Large Language Models as Optimizers (LLMaO). LLMs serve as optimizers within the algorithm to generate new solutions. This typically involves using the LLM in an iterative search process to enhance solution quality. Here, the algorithm and its parameters are usually crafted by humans.
One of the initial efforts to utilize LLMs as optimizers in algorithm design is by Yang et al. (2024). They leverage the in-context learning capabilities of LLMs to generate novel solutions for specific problems based on previously evaluated solutions. This method is applied iteratively to refine solutions further. Yang et al. (2024) have successfully demonstrated this technique across various domains, including continuous and combinatorial optimization, as well as machine learning tasks.
From an evolutionary algorithm (EA) perspective, using LLMs to generate solutions from existing data can be seen as analogous to search operators in EA. For instance, Liu et al. (2023a) introduce the use of LLMs as evolutionary operators to tackle multi-objective problems. This method involves breaking down a multi-objective problem into simpler single-objective tasks, with LLMs acting as black-box search operators for each sub-problem to suggest new solutions. In a related study, Liu et al. (2024c) explore the integration of LLMs within EAs, not just for generating solutions but also for guiding selection, crossover, and mutation processes. Meanwhile, Brahmachary et al. (2024) propose a new population-based evolutionary framework that includes both exploration and exploitation pools, with solutions being exchanged during the optimization process and LLMs generating solutions for both pools.
Differing from direct solution generation, Lange et al. (2024) investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their EvoLLM method, which shows superior performance over baseline algorithms in synthetic black-box optimization functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms could further improve the performance of EvoLLM. Moreover, Custode et al. (2024) present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations.
Beyond traditional optimization tasks, LLMaO has been widely adopted in prompt engineering for LLMs, a process often referred to as “automatic prompt optimization” (Zhou et al., 2022). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models (typically LLMs). Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity (Wang et al., 2023b), and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors (Guo et al., 2023), have been explored. Ma et al. (2024d) note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. To address these issues, they proposed a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models.
3.2. LLMs as Predictors (LLMaP)
LLMaP employs LLMs as surrogate models (Fig. 5) to predict the outcomes or responses of solutions, functioning either in a classification or regression context (Hao et al., 2024). Compared to other model-based predictors, such as the Gaussian process and conventional neural networks, 1) LLMs are capable of processing and generating human-like responses based on the training they have received on vast datasets. This capability allows them to understand and interpret complex patterns in data, making them suitable for tasks where traditional modeling techniques might struggle due to the intricacies in the data and complicated representations (Egami et al., 2024). 2) Pre-trained LLMs can significantly reduce the computational load and time required compared to training high-fidelity models (Jawahar et al., 2023; Hao et al., 2024).
[…]
Figure 5. Large Language Models as Predictors (LLMaP). LLMs are utilized iteratively in algorithms to predict a solution’s outcomes or responses, typically functioning in classification or regression tasks.
The majority of LLMaP works use LLMs as pre-trained models to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by Jawahar et al. (2023). It offers a cost-effective alternative for performance estimation in neural architecture search. Zhang et al. (2024e) introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, Li et al. (2023) introduce CodonBERT for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and was trained on over 10 million mRNA sequences from various organisms. Soares et al. (2024) demonstrate the use of transformer-based models in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (Alipour-Vaezi and Tsui, 2024) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (Zhang et al., 2024a).
For classification, Hao et al. (2024) introduce LAEA, which employs LLMs as surrogate models within evolutionary algorithms for both regression and classification, eliminating the need for costly model training. In another study, Chen et al. (2023b) develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance. Moving beyond binary classification, Bhambri et al. (2024) utilize LLMs to predict discrete actions for constructing reward shaping functions in Reinforcement Learning (RL). Their method demonstrated effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. Wang et al. (2024d) explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks. Lastly, Mehrdad et al. (2024) focus on automating the relevance judgment of query-item pairs in product searches. By finetuning LLMs, they achieved significant improvements, underscoring the adaptability and effectiveness of LLMs in complex search tasks.
LLMs can respond to more than numerical data, which are quite different from conventional prediction models. For example, Egami et al. (2024) utilize LLMs to generate imperfect annotation surrogates for documents in computational social science research, combined with high-quality labels to provide valid statistical inference for downstream analyses. In addition, the multi-modal LLMs demonstrate exceptional capability in utilizing multi-model information. An evolutionary strategy with a vision-language model has been developed by Wong et al. (2024b) for practical 3D car designs. The vision LLM is used to predict practicality degree for designs that satisfy the given text prompt.
3.3. LLMs as Extractors (LLMaE)
LLMaE employs LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then utilized in the enhancement of algorithm-based problem solving (Fig 6). This process leverages the unique capabilities of LLMs, such as text and code comprehension, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods (Wu et al., 2024c).
[…]
Figure 6. Large Language Models as Extractors (LLMaE). LLMs are employed to extract features or specific knowledge from target problem and/or algorithms to enhance problem-solving.
In the context of feature extraction, Kristiadi et al. (2024) explore the use of LLMs in accelerating Bayesian optimization in the molecular space. LLMs are used as fixed feature extractors for standard Bayesian optimization surrogate models, and parameter-efficient finetuning methods are leveraged to obtain the posterior of the LLM surrogate.
Beyond conventional feature extraction, LLMs excel in text comprehension and knowledge extraction. An example in LLM-assisted algorithm selection by Wu et al. (2024c) utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. In another innovative application, Du et al. (2024c) propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, Becker et al. (2024) discuss the use of LLMs in the automotive and supplier industries, specifically focusing on retrieval-augmented generation systems to improve information retrieval from technical documentation. Memduhoğlu et al. (2024) use LLM embeddings to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
The multi-modal capabilities of LLMs also distinguish them from traditional methods. Park et al. (2024) demonstrate that incorporating lexical information extracted from LLMs into an acoustic-based speaker diarization system through a probabilistic multi-modal decoding process and beam searches can significantly improve speech-processing performance.
3.4. LLMs as Designers (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 7). This utilization of LLMs extends their application beyond traditional boundaries, enabling them to actively participate in algorithm development by generating heuristics (Liu et al., 2024d), writing code snippets (Ma et al., 2024b), or formulating functions (Romera-Paredes et al., 2024) that can be integrated into algorithmic systems. By doing so, LLMs can significantly accelerate the algorithm design process, reduce human effort, and bring creativity and optimization to algorithm development (Liu et al., 2024d), which is difficult to achieve through conventional methods alone.
[…]
Figure 7. Large Language Models as Designers (LLMaD). LLMs are used to directly create algorithms or specific components, which are commonly incorporated iteratively to continuously search for better designs.
Function design is among the early applications of LLMaD. Eureka (Ma et al., 2024b) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for RL. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. Similarly, Auto MC-Reward (Li et al., 2024b) utilizes LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. Moreover, FunSearch (Romera-Paredes et al., 2024) adopts LLMs for function generation in an evolutionary framework with a multi-iland population management. FunSearch demonstrated promising results on both mathematical problems and combinatorial optimization problems.
EoH (Liu et al., 2024d), originates from AEL (Liu et al., 2023b), presents an early attempt to adopt LLM as a designer for automatic heuristic design. It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the algorithmic heuristic. EoH is applied to the key heuristic component design for combinatorial optimization problems, where it surpasses both traditional human-designed metaheuristics and deep-learning-based neural solvers. The application of EoH has expanded to include cost-aware acquisition function design for Bayesian optimization (Yao et al., 2024a), adversary attack algorithm design for image adversary attack (Guo et al., 2024b), and heuristic design for edge server task scheduling Yatong et al. (2024). LLaMEA (van Stein and Bäck, 2024) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (Ye et al., 2024a) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. With the reflection, better results on black-box cases (no problem-specific information is provided in the prompts) have been observed. Moreover, Yan et al. (2024) explore the use of LLMs to generate drafts of high-quality algorithms for electronic health records phenotyping, aiming to improve efficiency and scalability in observational health research and Hazra et al. (2024) combine LLMs with heuristic planning to generate cost-effective plans, showing superior performance in LLM planning approaches. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (Yao et al., 2024b) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
LLM-based agent design has also gained much attention. For example, ADAS (Hu et al., 2024) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents. They present a novel algorithm, meta agent search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains. Further studies on LLM-based agent systems are discussed in (Sun et al., 2024a) and (Liu et al., 2024e).
There has been a considerable effort on using LLMs for optimization modeling. Ahmed and Choudhury (2024) assess various LLMs for transforming linguistic descriptions into mathematical optimization problems, introducing a fine-tuning framework, LM4OPT, to enhance the performance of smaller models. Tang et al. (2024a) outline four key training dataset requirements for operational research LLMs and propose OR-Instruct, a semi-automated method for generating tailored synthetic data. Additionally, Mostajabdaveh et al. (2024) present a novel multi-agent framework that translates natural language into optimization models, utilizing relational identifier agents and a verification mechanism for improved assessment.
4. Search Methods
Incorporating LLMs within a search framework significantly enhances LLMs’ utility in algorithm design, making it the preferred option (Ma et al., 2024b; Liu et al., 2024d; Romera-Paredes et al., 2024). This section categorizes existing research according to the search methods employed, presents recent advancements, and discusses some of the limitations.
4.1. Sampling
The most straightforward search manner is by repeatedly instructing LLM to sample new designs (Zhou et al., 2022). The best sample is selected as the final design. However, recent studies have shown that simple sampling from LLMs can be costly (Zhang et al., 2024b).
4.1.1. Beam Search
Beam search explores multiple promising paths in the search space. For instance, in LLM-based prompt engineering (Zhou et al., 2022; Pryzant et al., 2023; Ma et al., 2024d), increasing the beam size significantly enhances performance (Zhou et al., 2022). Beyond prompt engineering, beam search is utilized in diverse applications: Text2Motion (Lin et al., 2023) employs it for robotic task planning, SayCanPay (Hazra et al., 2024) for action sequence planning in robots, enhancing speech-processing tasks, and in a confidence-aware method for structured data generation.
4.1.2. MCTS
Monte Carlo Tree Search (MCTS) is a tree search algorithm that uses random sampling to build a search tree and make decisions based on the results of these samples. It is particularly effective in problems with large and complex search spaces. For example, Dainese et al. (2024) propose GIF-MCTS, a code generation strategy using MCTS, in which nodes in the tree are programs and edges are actions. Each action taken from a parent node produces a new complete program. The upper confidence bound for trees is used to select which action to take. Similarly, VerMCTS (Brandfonbrener et al., 2024) constructs a search tree with progressive widening to effectively manage large action spaces defined by lines of code. In this search tree, expansion and evaluation are guided by the verifier, serving as a computationally inexpensive upper bound on the value function. Moreover, Wang et al. (2023b) regard prompt optimization as a strategic planning challenge, utilizing a principled planning algorithm based on MCTS to effectively explore the expert-level prompt space. The proposed PromptAgent generates precise insights and detailed instructions by analyzing model errors and providing constructive feedback.
4.2. Single-point-based Search
Single-point-based search methods iteratively refine a solution by leveraging neighborhood structures (Liu et al., 2024f) or specific search directions (Pryzant et al., 2023). While these methods can produce satisfactory results within a reasonable number of iterations, they may struggle with maintaining diversity and robustness in their search processes.
4.2.1. Hillclimb
Hillclimb search iteratively searches for better results, where the new one is generated based on the last one and the better one is survived. A basic version of Hillclimb is investigated by Zhang et al. (2024b) for algorithm design, where a heuristic is iteratively refined. The reasoning over the existing status is usually adopted to enhance each search step. For example, Li et al. (2024b) leverage LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards, which iteratively refines the reward function based on feedback from the agent’s interactions with the environment and Yang et al. (2023) develop a multi-module framework with innovative feedback mechanisms, demonstrating through both LLM-based and expert evaluations that LLMs can effectively produce scientific hypotheses that are both novel and reflective of reality.
4.2.2. Neighborhood Search
Neighborhood search methods explore the solution space by iteratively modifying a current solution to find an improved one in a structured neighborhood. Notably, the LLM-GS framework (Liu et al., 2024f) has introduced a scheduled neighborhood search method that leverages two distinct strategies: 1) Programmatic Neighborhood Generation, which involves generating a neighborhood of programs by selecting a node from the abstract syntax tree of a given program and substituting it with a subtree that is randomly generated following the production rules and sampling strategies of a domain-specific language; and 2) Latent Space Representation, where the neighborhood is defined in a latent space by training a variational autoencoder on a corpus of randomly generated DSL programs, thereby creating a more abstract and potentially more informative neighborhood structure. Additionally,  Jin et al. (2023) contribute to this line of research by defining neighborhood algorithms based on a minimum cosine similarity threshold of 60% between the embeddings of a query vector and code snippets, enhancing the search process by ensuring that algorithmic variations are semantically coherent and aligned with the underlying problem structure.
4.2.3. Gradient-based Search
Gradient-based search utilizes a (pseudo) gradient direction to generate new designs in each iteration. Unlike searches in a continuous space with a differentiable objective, calculating the gradient in the algorithm design space poses significant challenges. Consequently, a pseudo descent direction is often employed. Pryzant et al. (2023) introduce automatic prompt optimization, which improves prompts automatically by adjusting them based on natural language “gradients” derived from training data. Similarly, Tang et al. (2024c) introduce GPO, which draws parallels with gradient-based model optimizers for prompt optimization. Moreover, Nie et al. (2024) explore the use of LLMs as interactive optimizers for solving maximization problems in a text space through natural language and numerical feedback. For an accurate gradient, Guo et al. ([n. d.]) perform the gradient-based search in a mapped continuous space and proposes a collaborative optimization strategy that combines a gradient-based optimizer with an LLM for tackling complex non-convex optimization challenges. These works employ either pseudo gradients or gradients over mapped text space; however, a systematic study of gradients over general algorithm space is anticipated.
4.2.4. Reinforcement Learning
Reinforcement learning learns to make decisions by interacting with an environment to maximize rewards. GPTswarm (Zhuge et al., 2024) and Duan et al. (2023) leverage LLMs and RL techniques to optimize code, reducing complexity and enhancing efficiency. Similarly, Liu et al. (2024f) present a novel LLM-guided search framework (LLM-GS) aimed at addressing the sample inefficiency in state-of-the-art programmatic RL methods by utilizing the programming expertise of LLMs to boost search efficiency. Additionally, LLMs have been utilized in the design of RL systems, particularly in the creation of reward-shaping functions. Moreover, Bhambri et al. (2024) propose a framework called MEDIC to tackle the issue of sample inefficiency. They use LLMs to generate a guide policy for constructing reward-shaping functions.
4.3. Population-based Search
Population-based evolutionary search is the main tool investigated in LLM4AD papers due to its effectiveness and robustness in complex search space (O’Reilly and Hemberg, 2024; Wu et al., 2024a). The majority of these works use a simple genetic algorithm with greedy population management Liu et al. (2024d). Some of them investigate advanced population management including multi-island Romera-Paredes et al. (2024), quality-diversity Hu et al. (2024), and dominance-dissimilarity measurement Yao et al. (2024b). According to the number of objectives, they can be roughly categorized into single-objective search and multi-objective search.
4.3.1. Single-objective Evolutionary Search
Liu et al. (2024c) explore the use of LLMs as evolutionary operators in conventional EA, guiding them in solution selection, crossover, and mutation processes. Moreover, Brahmachary et al. (2024) propose a population-based evolutionary framework comprising exploration and exploitation pools, with LLMs facilitating solution generation for both pools and enabling communication between them during optimization. Lange et al. (2024) investigate the application of LLMs in designing evolution strategies, introducing a prompting strategy to enhance mean statistic performance.
On algorithm design tasks involving text and code search, the complicated search space poses challenges for diversity control and population management. The majority of workers adopt a greedy way (Liu et al., 2024d; Chen et al., 2024a; van Stein and Bäck, 2024; Li et al., 2024b; Narin, 2024; Guo et al., 2024b; Yao et al., 2024a). In these works, a population of individuals is maintained and only the ones with better fitness will survive. Romera-Paredes et al. (2024) adopt a multi-island evolutionary algorithm to explore a diverse population with a dynamically adjusted population size. Moreover, Wong et al. (2024a) use LLMs to create digital artifacts for creative and engineering applications, utilizing a multi-factorial evolutionary algorithm to drive an LLM. Quality diversity methods (Pugh et al., 2016) have also been investigated for this purpose, Lehman et al. (2023) and Hu et al. (2024) utilize the MAP-Elites algorithm with a grid of niches to generate diverse and high-quality solutions.
4.3.2. Multi-objective Evolutionary Search
Liu et al. (2023a) propose MOEA/D-LMO to use LLM to solve continuous multi-objective optimization problems. Benefiting from the decomposition-based framework, the in-context learning of LLM is used to generate new solutions for each subproblem. For multi-objective LLM instruction generation, Yang and Li (2023) develop InstOptima, which utilizes an LLM to simulate instruction operators and improve instruction quality to simultaneously optimize performance, length, and perplexity of the instruction. Moreover, Morris et al. (2024) introduce guided evolution for neural architecture design. It adopts NSGA-II and optimizes both accuracy and model size. In order to enhance the diversity control in multiobjective heuristic design, MEoH (Yao et al., 2024b) introduce a dominance-dissimilarity score for multi-objective heuristic design, achieving a good balance of exploration and exploitation in the heuristic search space.
4.4. Uncertainty-guided Search
Uncertainty-guided search harnesses the strengths of Bayesian Optimal Experiment Design (BOED) or Bayesian Optimization (BO) frameworks within the search process incorporating LLMs. It first establishes a prior that describes the initial belief about the parameters of interest and then employs uncertainty-driven strategies over the derived posterior to iteratively update the belief, ultimately enabling efficient decision-making during the search process in a sample-efficient manner.
This search paradigm with uncertainty has demonstrated its superiority in some scenarios. For example, Handa et al. (2024) introduce OPEN using LLMs to extract environment-relevant features and translate abstract feature queries into natural language questions to interact with users through conversations. Meanwhile, BOED is employed to identify the most informative query for efficiently learning preference information. To improve the multi-turn and decision-theoretic reasoning capabilities over a set of arbitrary items, Austin et al. (2024) formulate the natural language-based preference elicitation within a BO framework. The proposed PEBOL algorithm establishes a probabilistic prior over user preferences in the form of Beta distributions. In the domain of LLM decoding, Grosse et al. (2024) propose a probabilistic model named ULTS. This method frames the decoding process as a tree-search problem and aims to find the optimal path in the LLM search tree by sequential decision-making under uncertainty.
5. Prompt Strategies
(a) Pre-trained and domain LLMs
(b) Prompt engineering techniques
(c) LLMs
[…]
Figure 8. LLM types and promote engineering methods
Fig. 8(a) depicts the percentage of domain or pre-trained LLMs used in the literature. Among them, over 80% chose to use the pre-trained model without any specific finetuning, and about 10% fine-tuned the pre-trained model on domain datasets (Cummins et al., 2024) in which 4.4% are trained from scratch. Fig. 8(c) illustrates the top-used LLMs. LLM4AD papers show a strong preference for GPT models. GPT-4 and GPT-3.5 are the most-used LLMs, which in total take around 50%. Llama-2 is the most used open-source LLM. Once we have the pre-trained LLMs, prompt engineering is significant for effectively integrating LLMs in algorithm design. We introduce the main prompt engineering methods including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection (Sahoo et al., 2024) used in LLM4AD papers. Fig. 8(b) shows the number of papers involving different prompt engineering techniques.
5.1. Zero-Shot
Zero-Shot (ZS) prompting enables a model to comprehend and perform tasks without prior specific training on those tasks. In algorithm design, zero-shot prompting allows direct requests to the LLM for solutions or responses, examples including conversational replies (Sekulić et al., 2024), assertion design (Miftah et al., 2024), RL guidance (Zhang et al., 2024e), and molecule generation (Ye, 2024). The model leverages its pre-trained knowledge to generate these responses. Although this method is advantageous for swiftly producing solutions based on general knowledge, it may not consistently deliver tailored responses to the nuanced demands of complex algorithmic challenges (Zhang et al., 2024b).
5.2. Few-Shot
Few-Shot (FS) prompting involves providing the model with a handful of examples to illustrate the task before it attempts to solve a similar problem. This approach enhances the model’s understanding of the context and its ability to tailor responses to specific tasks. In algorithm design, this method proves particularly effective by presenting the model with examples of algorithms (Liu et al., 2023b), solutions (Yang et al., 2024), prompts (Wang et al., 2023b; Tang et al., 2024c), and codes (Ma et al., 2024a; Morris et al., 2024). The examples used in the prompt may be manually-designed seeds (Romera-Paredes et al., 2024; Hu et al., 2024) or derived from LLM-generated designs (Liu et al., 2024d; Wong et al., 2024a; Ma et al., 2024c; Morris et al., 2024; Wang et al., 2023b; Tang et al., 2024c; Custode et al., 2024). These samples are typically sorted to suggest a preference to enhance performance (Yang et al., 2024; Romera-Paredes et al., 2024).
5.3. Chain-of-Thought
Chain-of-Thought (CoT) prompting encourages the model to articulate intermediate steps or reasoning paths that lead to the final answer. This technique is particularly valuable in algorithm design, where understanding the step-by-step process or engaging in instructed reasoning over existing designs helps prevent outlier results and fosters effective algorithm development. For instance, Liu et al. (2024d) introduce several prompting strategies, one of which directs the LLM to reason over existing heuristics, summarize the general pipeline, and design a new one based on this reasoning. Custode et al. (2024) prompt the LLM to evaluate whether the current step size is sufficient for convergence and to suggest adjustments based on its reasoning. In addition, multiple LLM-based agents are employed within a CoT framework by Sun et al. (2024b) to enhance heuristic design for a boolean satisfiability problem solver.
5.4. Self-consistency
Self-consistency involves generating multiple answers or solutions to the same prompt and then synthesizing them to improve accuracy and reliability. In algorithm design, this could mean prompting the model multiple times for different approaches to solve a problem and then comparing these solutions to identify the most efficient or robust algorithm. This approach leverages the model’s ability to generate diverse solutions and builds a more comprehensive understanding of possible strategies. For example, Guan et al. (2023) set the number of responses as 10 and uses the majority vote to get the predicted label. As each response may provide different keywords or regular expressions, it takes the union of the keywords or regular expressions to create a candidate set. Moreover, Lehman et al. (2024) sample multiple revised codes from the same selected code and updates the population accordingly.
5.5. Reflection
Reflection in prompt engineering involves asking the model to critique or evaluate its own responses or solutions. After generating an algorithm/solution, the model can be prompted to reflect on the efficiency, potential flaws, or improvements. Ma et al. (2024d) investigate both implicit and explicit reflection in LLM-driven prompt optimization, which lets LLM analyze the errors and generate a reflection or feedback regarding the current prompt. An additional study by Ye et al. (2024b) incorporates short-term and long-term reflections in heuristic design, while Zhang et al. (2024c) focus on self-evaluation and reasoning over results from network detection. Furthermore, both Ma et al. (2024b) and Narin (2024) adopt LLM for RL reward function design with a reward reflection that monitors the scalar values of all reward components and the task fitness function at various policy checkpoints during training.
6. Applications
6.1. Optimization
In this subsection, we delve into the practical applications of LLMs in algorithm design for optimization. We categorize the existing literature into combinatorial optimization, continuous optimization, Bayesian optimization, prompt optimization, and optimization modeling based on the specific domains of optimization applications. Then we proceed to compare the various roles played by LLMs, the prompt strategies employed, and the specific problems or tasks to which they are applied. The comparative analysis is summarized in Table 1, where we list the names of the frameworks or methods proposed by the authors. For studies that do not explicitly name their methods, we assign appropriate designations in our article and denote them with asterisks for easy reference (e.g., MH-LLM* for (Sartori et al., 2024)).
6.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automatic algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (Liu et al., 2023b), ReEvo (Ye et al., 2024a) and EoH (Liu et al., 2024d). Differently, OPRO (Yang et al., 2024) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (Liu et al., 2024b) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (Huang et al., 2024c) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance.
In addition to routing problems, other combinatorial optimization problems that have been investigated include cap set (Romera-Paredes et al., 2024), bin packing (Liu et al., 2024d), flow job shop scheduling (Liu et al., 2024d) and social networks problems (Sartori et al., 2024).
6.1.2. Continuous Optimization
In the realm of continuous single-objective optimization, LLaMEA (van Stein and Bäck, 2024) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (Hansen and Ros, 2010) within IOHexperimenter benchmarking tool (de Nobel et al., 2024), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (Lange et al., 2024) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (Yang et al., 2024) illustrates that LLMs can effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (Brahmachary et al., 2024) devises an explore-exploit policy using LLMs for solution generation, the method has been tested and validated in both benchmark functions as well as industrial engineering problems. Different with directly employing LLMs for generating solutions, LAEA (Hao et al., 2024) introduces LLM-based surrogate models for regression and classification tasks to assess the quality of new solutions for selection. This approach has been validated on 2D test functions using nine mainstream LLMs.
Multi-objective optimization problems (MOPs) seek to identify a set of optimal solutions, referred to as Pareto optimal solution set. An initial exploration of utilizing LLMs to tackle MOPs is introduced in MOEA/D-LMO (Liu et al., 2023a). Benefiting from the decomposition-based framework, the in-context learning process of LLMs is easily incorporated to generate candidate solutions for each subproblem derived from the original MOP. In the realm of large-scale MOPs, LLM-MOEA* (Singh and Bali, 2024) showcases the inferential capabilities of LLMs in multi-objective sustainable infrastructure planning problem. The study highlights the LLM’s proficiency in filtering crucial decision variables, automatically analyzing the Pareto front, and providing customized inferences based on varying levels of expertise. Additionally, CMOEA-LLM (Wang et al., 2024b) leverages LLMs with evolutionary search operators to address the challenges of constrained MOPs and exhibits robust competitiveness in DAS-CMOP test suite (Fan et al., 2020).
6.1.3. Bayesian Optimization
Bayesian optimization (BO) is a model-based optimization paradigm for solving expensive optimization problems and has found wide application in various real-world scenarios (Frazier and Wang, 2016). It typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to select potential solutions in a sample-efficient manner. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (Zhang et al., 2023) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (Liu et al., 2024a) incorporates LLM capabilities to enhance the efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from Bayesmark (Uber, 2020) and HPOBench (Eggensperger et al., 2021). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (Ramos et al., 2023) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (Dinh et al., 2022) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (Yao et al., 2024a) introduces a novel paradigm to design AFs automatically for cost-aware BO, the approach showcases remarkable efficiency and generalization for synthetic functions and HPO tasks on JAHS-Bench-201 (Bansal et al., 2022). FunBO (Aglietti et al., 2024) discoveres novel and well-performing AFs for BO by extending FunSearch (Romera-Paredes et al., 2024), the discovered AFs are evaluated on various global optimization benchmarks in and out of the training distribution and HPO tasks for RBF-based SVM and AdaBoost algorithms.
6.1.4. Prompt Optimization
Prompt optimization aims to identify the most effective task prompt that maximizes the performance of the LLM on a specific task dataset. Despite of requiring specialized training for each specific task, traditional discrete or continuous approaches (Prasad et al., 2022; Li and Liang, 2021) typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (Zhou et al., 2022) utilizes the LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (Honovich et al., 2022) and Big-Bench Hard (BBH) (Suzgun et al., 2022). OPRO (Yang et al., 2024) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on GSM8K (Cobbe et al., 2021) and BBH. Inspired by the numerical gradient descent method, APO (Pryzant et al., 2023) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. Similar practices are also found in the gradient-inspired LLM-based prompt optimizer named GPO (Tang et al., 2024c), as well as the collaborative optimization approach (Guo et al., 2024a) integrating a gradient-based optimizer and an LLM-based optimizer. Differently, Guo et al. (2023) introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts, harnessing the benefits of evolutionary algorithms that strike a good balance between exploration and exploitation.
6.1.5. Optimization Modeling
Optimization modeling (Berry and Houston, 1995) aims to construct a mathematical model of a real-world optimization problem in a standard form, represented with an objective function and a set of constraints (AhmadiTeshnizi et al., 2023; Tang et al., 2024a). In addition to resulting significant differences in solving efficiency, creating an efficient formulation often demands expertise in optimization skills and specialized domain knowledge in various applications. Recently, LLMs have opened up promising avenues for automating optimization modeling, facilitating the formulation of problems and the implementation of solutions, thereby democratizing expert knowledge. AhmadiTeshnizi et al. (2023) propose OptiMUS, an LLM-based agent designed to formulate and solve optimization problems interpreted from natural language. The performance of the method is evaluated on a proposed novel dataset named NLP4LP (AhmadiTeshnizi et al., 2023), which includes 41 linear programming and 11 mixed integer linear programming problems. Despite its outstanding performance, the heavy reliance on sensitive data submission to proprietary LLMs can pose data privacy concerns in industrial applications.
To solve this issue, Tang et al. (2024a) propose training open-source LLMs for optimization modeling that are applicable to real-world industries. They develop and implement OR-Instruct, a semi-automated process for generating synthetic data customized to meet specific requirements. The best-performing resulting model, ORLM, demonstrates state-of-the-art performance on NL4Opt (Ramamonjison et al., 2023), MAMO (Huang et al., 2024a), and the proposed IndustryOR benchmark.
Ahmed and Choudhury (2024) develop a progressive fine-tuning framework named LM4OPT to enhance LLMs’ specificity for optimization problem formulation task on NL4Opt dataset. Wasserkrug et al. (2024) develop Decision Optimization CoPilot (DOCP), which is accessible to a broad range of decision-makers lacking problem-dependent modeling skills. This tool aids in decision-making by interacting with the user in natural language and utilizes human’s knowledge and feedback to create and solve a problem-specific optimization model. The experiments are conducted targeting three presented requirements for DOCP on facility location problems.
6.1.6. Other Applications
In addition to optimization applications that can be categorized into various domains as mentioned above, we have also discovered applications of LLMs in other optimization-related fields, which may intersect to varying degrees with the aforementioned classifications. For ease of reference, we have categorized them as ”other applications” and will introduce them separately below.
Wu et al. (2024c) propose AS-LLM, which harnesses the representation capabilities of LLMs to extract algorithm features for algorithm selection tasks. The approach integrates a feature selection module to identify critical features and leverages the similarity between algorithm and problem representations. In addition to demonstrating robust performance advantages on the ASlib (Algorithm Selection Library) benchmark (Bischl et al., 2016) in experimental studies, the authors also offer a rigorous upper bound on the model complexity of AS-LLM. Ma et al. (2024a) introduce LLaMoCo, the first instruction-tuning framework tailored to adapt general-purpose LLMs for optimizing code generation. Benefiting from a well-formatted instruction set and a novel two-phase training strategy, the fine-tuned LLMs exhibit robust and superior optimization performance, reducing the requirement for extensive domain knowledge and effort in prompt design. StrategyLLM (Gao et al., 2023) employs four LLM-based agents - the strategy generator, executor, optimizer, and evaluator - in a collaborative manner. The approach enables LLMs to reason inductively and deductively, resulting in the formulation of more generalizable and consistent few-shot prompts compared to CoT prompting methods. Experimental results demonstrate the effectiveness and reliability of StrategyLLM across four challenging tasks, encompassing math reasoning, commonsense reasoning, algorithmic reasoning, and symbolic reasoning.
Table 1. An Overview of Optimization Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Prompt Strategy Specific Problems or Tasks
Combinatorial Optimization AEL (Liu et al., 2023b) LLMaD FS, CoT TSP
ReEvo (Ye et al., 2024a) LLMaD CoT, RF TSP
OPRO (Yang et al., 2024) LLMaO FS TSP
LMEA (Liu et al., 2024b) Mixed FS TSP
MLLM (Huang et al., 2024c) Mixed ZS, FS CVRP
FunSearch (Romera-Paredes et al., 2024) LLMaD FS Cap Set Problem, Online BPP
EoH (Liu et al., 2024d) LLMaD FS, OS, CoT TSP, Online BPP, FSSP
MH-LLM* (Sartori et al., 2024) LLMaD ZS, FS Social Networks Problem
Continuous Optimization LLaMEA (van Stein and Bäck, 2024) LLMaD CoT BBOB
EvoLLM (Lange et al., 2024) LLMaO FS BBOB, Neuroevolution
OPRO (Yang et al., 2024) LLMaO FS Linear Regression
LEO (Brahmachary et al., 2024) LLMaO FS Numerical Benchmarks, Industrial Engineering Problems
LAEA (Hao et al., 2024) LLMaP FS, CoT Ellipsoid, Rosenbrock, Ackley, Griewank
MOEA/D-LMO (Liu et al., 2023a) LLMaO FS ZDT, UF
LLM-MOEA* (Singh and Bali, 2024) LLMaE ZS Multi-objective Sustainable Infrastructure Planning Problem
CMOEA-LLM (Wang et al., 2024b) LLMaO FS DAS-CMOP
Bayesian Optimization HPO-LLM* (Zhang et al., 2023) LLMaO FS HPOBench
LLAMBO (Liu et al., 2024a) Mixed FS, ZS, CoT Bayesmark, HPOBench
BO-LIFT* (Ramos et al., 2023) LLMaD FS Catalyst Optimization
EvolCAF (Yao et al., 2024a) LLMaD FS, OS, CoT Synthetic Functions, JAHS-Bench-201
FunBO (Aglietti et al., 2024) LLMaD FS Synthetic Functions, HPO
Prompt Optimization APE (Zhou et al., 2022) LLMaO FS Instruction Induction, BBH
OPRO (Yang et al., 2024) LLMaO FS GSM8K, BBH, MultiArith, AQuA
APO (Pryzant et al., 2023) LLMaO ZS, FS NLP Benchmark Classification Tasks
GPO (Tang et al., 2024c) LLMaO FS Reasoning, Knowledge-intensive, NLP Tasks
MaaO (Guo et al., 2024a) LLMaO FS NLU Tasks, Image Classification Tasks
EvoPrompt (Guo et al., 2023) LLMaO CoT, FS Language Understanding and Generation Tasks, BBH
Optimization Modeling DOCP (Wasserkrug et al., 2024) LLMaO FS, CoT Facility Location Problems
OptiMUS (AhmadiTeshnizi et al., 2023) Mixed ZS, FS NL4LP
ORLMs (Tang et al., 2024a) LLMaD ZS NL4Opt, MAMO, IndustryOR
LM4OPT (Ahmed and Choudhury, 2024) LLMaD FS, ZS NL4Opt
Other Applications AS-LLM (Wu et al., 2024c) LLMaE ZS Algorithm Selection
LLaMoCo (Ma et al., 2024a) LLMaD FS Optimization Code Generation
StrategyLLM (Gao et al., 2023) Mixed ZS, FS, CoT Math & Commonsense & Algorithmic & Symbolic Reasoning
6.2. Machine Learning
Table 2. An Overview of Machine Learning Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Prompt Strategy Specific Problems or Tasks
Task Planning Zero-Planner (Huang et al., 2022) LLMaP ZS, FS VirtualHome Tasks
SayCan (Ahn et al., 2022) LLMaP FS Real-world Kitchen Tasks with Robots
LLM-GM (Huang et al., 2023) LLMaP FS Real-world Kitchen Tasks with Robots
Text2Motion (Lin et al., 2023) LLMaP FS Long Sequence Tasks for Robots
ProgPrompt (Singh et al., 2023) LLMaD FS, COT VirtualHome Tasks
SayCanPay (Hazra et al., 2024) LLMaP FS VirtualHome Tasks, Ravens Tasks, BabyAI Tasks
Reinforcement Learning LFG (Shah et al., 2023) LLMaP FS, CoT ObjectNav Tasks, Real-world Tasks
SLINVIT (Zhang et al., 2024e) LLMaP ZS, FS ALFWorld Tasks, InterCode Tasks, BlocksWorld Tasks
MEDIC (Bhambri et al., 2024) LLMaP ZS BabyAI Tasks
Eureka (Ma et al., 2024b) LLMaD FS, CoT IsaacGym Tasks, Bidexterous Manipulation Tasks
EROM (Narin, 2024) LLMaD ZS, CoT IsaacGym Tasks
LLM-MOE (Du et al., 2024c) LLMaP FS, CoT Intelligent Networks
Neural Architecture Search EvoPrompting (Chen et al., 2023a) LLMaD FS, ZS, CoT MNIST Dataset, CLRS Algorithmic Reasoning
HS-NAS (Jawahar et al., 2023) LLMaP FS, ZS, CoT Machine Translation Tasks
LLMatic (Nasir et al., 2023) LLMaD FS, ZS, CoT CIFAR-10 Dataset, NAS-bench-201 Benchmarks
LAPT (Zhou et al., 2024a) LLMaD ZS, FS NAS201, Trans101, DARTs
LLM-GE (Morris et al., 2024) LLMaD FS, ZS, CoT CIFAR-10 Dataset
Graph Learning LLM-Critical (Mao et al., 2024) LLMaD FS, ZS, CoT Critical Node Identification
LLM-GNN (Chen et al., 2023b) LLMaP FS, CoT Label-free Node Classification
ReStruct (Chen et al., 2024c) LLMaE FS, ZS Meta-structure Discovery
AutoAlign (Zhang et al., 2024d) LLMaE FS, ZS Entity Type Inference
KSL (Feng et al., 2023b) LLMaP FS Knowledge Search
Dataset Labeling Inter-Classier (Chiquier et al., 2024) LLMaO FS, ZS iNaturalist Datasets, KikiBouba datasets
DataSculpt (Guan et al., 2023) LLMaP FS Label Function Design
Other Applications LLM2FEA (Wong et al., 2024a) LLMaP FS, CoT Objective-oriented Generation
tnGPS (Zeng et al., 2024) LLMaD FS, OS, CoT Tensor Network Structure Search
L-AutoDA (Guo et al., 2024b) LLMaD FS, OS, CoT Adversarial Attack
In this subsection, we investigate the applications of LLMs in the machine learning domain, focusing on their contribution to algorithmic design. These applications are summarized in Table 2.
6.2.1. Task Planning with LLMs
The advent of LLMs has garnered significant interest in their application to planning tasks, owing to their remarkable generative capabilities of instructions in the form of natural language. Pioneering work by Huang et al. (2022) demonstrates the effectiveness of LLMs in leveraging world knowledge to produce actionable plans for VirtualHome environments. This method was further enhanced by Ahn et al. (2022), who employed a value function to ground the world knowledge, enabling the derivation of an optimal policy for robotic task execution in lifelike kitchen settings. In subsequent research, Huang et al. (2023) broaden the scope of grounded information in a more adaptive way. Moreover, Lin et al. (2023) utilize greedy search in combination with heuristic algorithms to refine the planning system. Distinctly, Singh et al. (2023) tackle planning challenges by integrating programming languages, thereby generating code to streamline planning processes. A more comprehensive approach, named SayCanPay (Hazra et al., 2024), is developed for planning by considering both the affordance function and reward within the task framework.
6.2.2. Reinforcement Learning
Reinforcement Learning (RL) has been the de facto standard for sequential decision-making tasks, and recently, the synergy between RL and LLMs has emerged as a novel trend in the domain. This convergence mirrors the dynamics of task planning, yet places RL at the core of its methodology. Many of the LLM4AD papers on RL is for automatically designing the reward functions (Bhambri et al., 2024; Ma et al., 2024b; Narin, 2024). In addition, Shah et al. (2023) investigate the employment of LLMs for heuristic planning to steer the search process within RL frameworks. Zhang et al. (2024e) integrate LLMs into RL by introducing a Kullback-Leibler divergence regularization term that aligns LLM-driven policies with RL-derived policies. LLMs have also extended their reach to multi-agent RL scenarios, as shown by Du et al. (2024c), who illustrates their application within a Mixture-of-Experts system to direct RL models in the realm of intelligent network solutions.
6.2.3. Neural Architecture Search
Neural Architecture Search (NAS), which is a significant focus within the AutoML community, has been investigated in many LLM4AD papers. For example, Chen et al. (2024a) have integrated LLMs with evolutionary search to successfully generate NAS code for diverse tasks. Nasir et al. (2023) introduce a quality-diversity algorithm tailored to NAS, producing architectures for CIFAR-10 and NAS-bench-201 benchmarks. Moreover, Morris et al. (2024) introduce guided evolution for the development of neural architectures and suggest the concept of the evolution of thought in algorithm design. Except for using LLM for design, Jawahar et al. (2023) employ LLMs in predicting NAS performance, combining this approach with evolutionary search to effectively create novel network architectures. In contrast to the LLM-based architecture design and performance prediction, Zhou et al. (2024a) explore the adoption of LLMs for transferring design principles to narrow and guide the search space.
6.2.4. Graph Learning
Graph learning is another application with the advancing capabilities of LLMs in symbolic reasoning and graph processing. For example, Chen et al. (2023b) apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both Mao et al. (2024) and Chen et al. (2024c) adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. Zhang et al. (2024d) introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and Feng et al. (2023b) develop the knowledge search language to effectively conduct searches within knowledge graphs.
6.2.5. Dataset Labeling
LLMs have been used for mining semantic and multi-modal information from datasets. LLMs are employed to train interpretable classifiers to extract attributes from images (Chiquier et al., 2024) and to generate label functions for weakly supervised learning (Guan et al., 2023).
6.2.6. Other Applications
Other Applications of LLMs extend to a myriad of machine learning tasks. Notably, Wong et al. (2024a) capitalize on multi-task learning and the iterative refinement of prompts to foster innovative design approaches. Zeng et al. (2024) integrate LLMs with evolutionary search to develop a heuristic function aimed at efficiently sifting through candidate tensor sets representing the primal tensor network. Furthermore, Guo et al. (2024b) have blazed a trail in employing LLMs to generate novel decision-based adversarial attack algorithms, thus opening up a new diagram for the automatic assessment of model robustness.
6.3. Science Discovery
This subsection is dedicated to exploring scientific discoveries related to algorithm designs. Table. 3 summarizes the related works in this domain. For an in-depth discussion and review of LLMs in scientific discovery, readers are referred to survey papers (Birhane et al., 2023; Hadi et al., 2024; Jablonka et al., 2023).
6.3.1. General Science Discovery
In the realm of general scientific discovery, several frameworks have been developed that leverage LLMs. Notably, Du et al. (2024b) introduce LLM4ED, a framework that employs iterative strategies, including a black-box optimizer and evolutionary operators, to generate and optimize equations. This approach has shown significant advancements in stability and usability for uncovering physical laws from nonlinear dynamic systems. Similarly, Shojaee et al. (2024) present LLM-SR, which combines LLMs’ extensive scientific knowledge and code generation capabilities with evolutionary search. This framework excels in proposing and refining initial equation structures based on physical understanding, outperforming traditional symbolic regression methods in discovering physically accurate equations across multiple scientific domains. Another innovative framework, SGA, introduced by Ma et al. (2024c), is a bilevel optimization framework that merges the abstract reasoning capabilities of LLMs with the computational power of simulations. This integration facilitates hypothesis generation and discrete reasoning with simulations for experimental feedback and optimization of continuous variables, leading to improved performance in discovering constitutive laws and designing molecules.
6.3.2. Chemistry, Biology & Physics
In the field of chemistry, LLMs can be applied not only to conventional molecular generation and design (Ishida et al., 2024; Bhattacharya et al., 2024), but also to specialized areas such as drug molecule design (Ye, 2024), chemical reaction prediction and optimization (Ranković and Schwaller, 2023), and catalyst design (Wang et al., 2024a), providing customized solutions. Furthermore, LLMs have also shown promising application prospects in materials discovery, synthesis route planning, green chemistry, and other areas. These studies not only demonstrate the advantages of language models in molecular representation and optimization but also provide new tools and ideas for chemical research.
In biology, LLMs are increasingly being used for tasks such as protein engineering (Lv et al., 2024; Shen et al., 2024), enzyme design (Teukam et al., 2024), and biological sequence analysis (Feng et al., 2024). By combining LLMs with vast amounts of biological data, they can more accurately predict interactions between biological molecules and significantly improve the efficiency of bioinformatics workflows. This has important implications for drug discovery and therapeutic protein design. The unique sequence generation and optimization capabilities of LLMs offer new possibilities for tackling combinatorial optimization challenges in biomacromolecular design.
Although applications in physics are relatively fewer, some emerging work has demonstrated the broad prospects of LLMs. Pan et al. (2024) use multi-step prompt templates to prove that LLMs can perform complex analytical calculations in theoretical physics. Quantum computing algorithm design, physics simulation optimization, and computational methods in condensed matter.
6.3.3. Mechanics
MechAgents (Ni and Buehler, 2024) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, Du et al. (2024a) adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to state-of-the-art models. Another example in fluid dynamics has been investigated by Zhu et al. (2024), which introduces FLUID-LLM, a novel framework that integrates pre-trained LLMs with spatiotemporal-aware encoding to enhance the prediction of unsteady fluid dynamics, addressing the challenge of complex geometries in computational fluid dynamics. Furthermore, Buehler (2023) study LLM-based methods for forward and inverse mechanics problems including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.
Table 3. An Overview of Science Discovery Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Specific Problems or Tasks
General Scientific Equation Discovery Bilevel (Ma et al., 2024c) LLMaD Physical Scientific Discovery
LLM-SR (Shojaee et al., 2024) LLMaD Scientific Equation Discovery
LLM4ED (Du et al., 2024b) LLMaD Equation Discovery
Chemical ChatChemTS (Ishida et al., 2024) LLMaD Molecule Design
Debjyoti Bhattacharya, et al. (Bhattacharya et al., 2024) LLMaO Molecule Design
Agustinus Kristiadi, et al. (Kristiadi et al., 2024) LLMaE Molecule Design
BoChemian (Ranković and Schwaller, 2023) LLMaE Chemical Reaction
Multi-modal MoLFormer (Soares et al., 2024) LLMaP Battery Electrolytes Formulation Design
CataLM (Wang et al., 2024a) LLMaP Catalyst Design
Gavin Ye (Ye, 2024) LLMaD Drug Design
DrugAssist (Ye et al., 2023) LLMaO Drug Design
Biology MLDE (Tran and Hy, 2024) LLMaP Protein Design
X-LoRA (Buehler and Buehler, 2024) LLMaP Protein Design
CodonBERT (Li et al., 2023) LLMaP mRNA Design and Optimization
Revisiting-PLMs (Hu et al., 2022b) LLMaP Protein Function Prediction
Mechanics FLUID-LLM (Zhu et al., 2024) LLMaP Computational Fluid Dynamics
MechAgents (Ni and Buehler, 2024) LLMaD Mechanics Design
MeLM (Buehler, 2023) LLMaP, LLMaD Carbon Nanotubes and Proteins Design
Mengge Du, et al. (Du et al., 2024a) LLMaD Nonlinear Dynamics Equation Discovery
6.4. Industry
This subsection explores the transformative impact of LLMs on algorithm design across various industries. As we envision the future of manufacturing, LLMs are poised to play a pivotal role. Existing developments focus on enhancing research and investment to refine these models while assessing their performance in terms of safety, bias, and utility. LLMs can be seen as a gateway between humans and machines, facilitating informed decision-making.
For instance, LLMs are instrumental in building a comprehensive intelligent 6G network system (Long et al., 2024), addressing challenges such as low latency, high-frequency bands, high bandwidth, high transmission rates, and overall intelligence (Zhou et al., 2024b). Moreover, LLMs have the potential to revolutionize the telecom industry by streamlining tasks and reducing the need for extensive manpower and engineering expertise. However, it is crucial to address existing challenges to fully realize their potential (Maatouk et al., 2024).
Recently, the application of LLMs in Electronic Design Automation (EDA) has emerged as a promising area of exploration. LLM-based solutions have been developed to enhance interactions with EDA tools (Fang et al., 2024). In particular, within the VLSI design flow, it is essential to verify that the implementation conforms to its specifications. To achieve this, system verilog assertions need to be generated automatically. Generative AI techniques, such as LLMs, can be effectively utilized to facilitate this process (Miftah et al., 2024).
To enhance the reliability and availability of cloud services, conducting root cause analysis (RCA) for cloud incidents is essential. RCACopilot (Chen et al., 2024b) effectively matches incoming incidents with the appropriate incident handlers based on their alert types. It aggregates critical runtime diagnostic information, predicts the incident’s root cause category, and provides explanatory results without the need for manual investigations.
More applicable scenarios include various industrial design challenges, such as chip design (Chang et al., 2024), car shape design (Wong et al., 2024b), aircraft concept design (Lovaco et al., 2024) UI design (Feng et al., 2023a), and robot design (Yu et al., 2023). LLMs primarily leverage billions of web resources. However, practical applications require fine-tuning with thousands of specific representations of design intent to recognize the distinct patterns associated with each design task.
Other traditional industrial tasks, such as itinerary planning, can also be enhanced by leveraging the power of LLMs (Xie et al., 2024). LLM-based open-domain urban itinerary planning (Tang et al., 2024b) combines spatial optimization with LLMs to create detailed urban itineraries tailored to users’ requests. This approach extends the scope of itinerary planning by generating personalized urban itineraries based on user requests in natural language.
7. Challenges of LLMs in Algorithm Design
7.1. Scalability
One of the primary limitations of LLMs in algorithm design is their scalability, particularly for handling long inputs and outputs. LLMs are limited by a fixed context size, restricting the amount of information they can process at once, which is a challenge for complex algorithmic tasks requiring detailed specifications and extensive content. Additionally, even when input and output lengths are sufficient, LLMs struggle with comprehension and reasoning over long inputs, impacting their effectiveness in designing complex algorithms. For instance, LLMs tend to solve only low-dimensional problems when used as optimizers (Yang et al., 2024), and they are employed to design a function or heuristic component of an algorithm rather than the entire algorithm when used as designers (Liu et al., 2024d).
7.2. Interpretability
Interpretability is another challenge when using pre-trained LLMs for algorithm design. It is difficult to trace how specific inputs lead to particular outputs in LLMs due to their black-box nature. This opacity can be problematic, especially in critical applications, e.g., some industry applications, where understanding the decision-making process is essential for trust and reliability. Efforts to enhance the interpretability involve techniques such as feature visualization, model simplification (Liu et al., 2023a), and the study of the contribution of different parts of the data to the model’s behavior (Allen-Zhu and Li, 2023). Despite these efforts, achieving principled interpretability without compromising the performance of LLMs remains a key area of ongoing research.
7.3. Security
Integrating LLMs into algorithm design introduces significant security risks, including the mishandling of personal data and the potential creation of harmful code. LLMs are trained on extensive datasets that often contain sensitive information, posing a risk of privacy breaches if these models inadvertently generate outputs influenced by this data. Additionally, the autonomous nature of LLMs can lead to the unintentional development of flawed or malicious algorithms, a critical concern in sectors like financial services and personal data management (Zhao et al., 2024). To address these issues, it is vital to establish stringent data governance and model training protocols that prevent the inclusion of sensitive data, rigorously test algorithms under various conditions to prevent unintended behaviors, and continuously update security measures to counteract emerging vulnerabilities.
7.4. Cost
The utilization of LLMs in algorithm design also faces challenges related to costs, both in terms of model training and usage. Training domain LLMs requires significant computational resources, often involving extensive GPU or TPU usage over several weeks or even months, which incurs high financial and time costs. Once trained, the inference cost—the computational expense involved in applying the model to algorithm design—also becomes a critical factor, especially when the model is used frequently or requires rapid responses, such as the evolutionary search used in many LLM4AD works (Romera-Paredes et al., 2024). For practical applications in algorithm design, where multiple iterations and tests might be necessary, these costs can accumulate, making the use of LLMs potentially less viable for researchers without substantial computational budgets.
7.5. Idea Innovation
Despite the numerous works on LLM4AD, their ability to produce truly novel ideas is still under scrutiny. LLMs excel at interpolation within the range of their training data but may struggle with extrapolation to new, unseen scenarios. In the context of algorithm design, this means that while LLMs can effectively suggest algorithms or solutions that closely resemble or incrementally improve upon existing ones, they may not be as effective at pioneering radically new algorithmic concepts or methods. The reliance on large volumes of training data also raises concerns about the model’s dependency on historical data, which may bias its outputs towards conventional methods and potentially stifle innovation. While LLM does show the capability of generating new ideas in some tasks (Yao et al., 2024a), how to continuously explore practical novel ideas remains a challenge (Hu et al., 2024).
8. Future Directions
8.1. Domain LLM for Algorithm Design
Instead of using a general pre-trained general-purpose LLM, it is worthwhile studying how to train an LLM specifically for automatic algorithm design tasks. The following aspects can be explored in developing domain LLMs: 1) Training domain LLMs are costly and resource-consuming. On the one hand, advanced light fine-tuning techniques (e.g., LoRA (Hu et al., 2022a)) could be used for efficient domain adaptation. On the other hand, the size of algorithm LLMs for a specific application can be reduced with the help of domain data and knowledge. 2) Generating and collecting domain data for algorithm design poses challenges. Unlike general code generation or linguistic process tasks, there is no large and formatted data specific for algorithm design. Some Github repositories on algorithms might be helpful, but the resources should be scanned and cleaned for algorithm design. Lehman et al. (2024) provide an example of generation domain data for LLMs, where the samples generated during searching are used as the training dataset to fine-tune a pre-trained model resulting in better performance on the target problem. 3) Instead of learning a text and code generation model, how to learn the algorithm development ideas and the algorithmic reasoning ability remain unexplored (Wang and Zhou, 2024; OpenAI, 2024).
8.2. Multi-modal LLM for Algorithm Design
The primary focus of existing LLM4AD works is on utilizing the text comprehension and generation capabilities of LLMs, whether in language, code, or statistics. One advantage of LLMs compared to traditional model-based optimization is their ability to process multi-modal information like humans, which is rarely studied. Some attempts have been conducted to showcase the advantages of incorporating multi-modal information in algorithm design, as evidenced by Huang et al. (2024c) and Narin (2024). It is anticipated that more methods and applications utilizing multi-modal LLMs will be developed in the future.
8.3. Interaction with Human Experts
Further research is needed to explore the interaction between LLMs and human experts in algorithm design (Li et al., 2024a). For example, in LLMaD works, LLMs can be seen as intelligent agents, making it feasible for human experts to step in and take over tasks such as generating, modifying, and evaluating algorithms. Investigating how to facilitate efficient and productive collaboration between LLMs and human experts would be valuable. Ideas and techniques in collective intelligence (Malone and Bernstein, 2022) can be used for this purpose.
8.4. LLM-based Algorithm Assessment
LLMs can be helpful in algorithm assessment. Some attempts have been carried out to automatically evaluate algorithms and assess algorithm design. For example, Sterbini and Temperini (2024) show that LLMs can be used to analyze algorithm descriptions and compare them to the task description given. Such text similarity correlates well both with the grades given by the teacher and with the grades given by the students. OMNI-EPIC (Faldor et al., 2024) utilizes foundation models to automatically produce code that defines the next engaging and learnable tasks. By generating environments and reward functions, OMNI-EPIC has the capability to create a wide range of simulated learning tasks for algorithm evaluation. We expect more studies on LLM-based algorithm assessment.
8.5. Understand the Behavior of LLM
In the majority of works, LLM works as a black-box model. The interpretation of LLM’s behavior not only enriches our understanding of LLM’s behavior but also benefits cases where it is challenging or costly to directly request LLMs. Some attempts have been made to approximate and understand the in-context learning behavior of LLM in solution generation. For example, Liu et al. (2023a) has designed a white-box linear operator to approximate the results of LLM in multi-objective evolutionary optimization. In spite of these preliminary attempts, how to interpret LLM’s behavior is still an open question in many algorithm design cases including heuristic generation and idea exploration.
8.6. Fully Automatic Algorithm Design
Fully automatic algorithm design faces two primary challenges: 1) the generation of novel algorithmic ideas and 2) the creation of complex, lengthy code. While the generation of new ideas has been investigated in some works (Liu et al., 2024d), the complete design of algorithms (instead of a heuristic component), including both heuristic components and detailed code implementations, remains a challenge. Existing applications typically focus on automating components within predefined algorithm frameworks rather than creating new algorithms from scratch. Future research needs to tackle these complexities to advance the field of fully automatic algorithm design.
8.7. Benchmarking LLM4AD
Benchmarking allows for a fair, standardized, and convenient comparison, which can identify best practices and enable generating new insights for innovation. While we are glad to witness the emergence of diverse research works and applications, LLM4AD still lacks a benchmarking on either algorithm design test sets for systematic and scientific assessment or a strict pipeline and evaluation criteria on large language models for algorithm design. Several attempts have been made on related topics such as mathematical reasoning (Lin et al., 2024) and planning (Valmeekam et al., 2024). (Veličković et al., 2022) presents an algorithm test set built on some basic algorithms and (McLeish et al., 2024) demonstrates the superior performance of LLM when compared to neural algorithm reasoners. In the future, more benchmarks are expected and they are going to play a crucial role in advancing LLM4AD.
9. Conclusion
This paper has provided a systematic and up-to-date survey on large language models for algorithm design (LLM4AD). By systematically reviewing a significant corpus of main contributions in this emerging research field, this paper not only highlights the current state and evolution of LLM applications in algorithm design but also introduces a novel multi-dimensional taxonomy with comprehensive lists of papers that categorize the roles and functionalities of LLMs in this domain. This taxonomy serves as a framework for both academic and industrial researchers to understand and navigate the complex landscape of algorithm design using LLMs. We have also discussed the limitations and challenges currently faced by the field, such as issues of scalability, interpretability, and security. Moreover, the proposed future research directions aim to inspire and guide subsequent studies.
As we look forward, it is clear that the intersection of LLMs and algorithm design holds promising potential for revolutionizing how algorithms are developed and implemented. The ability of LLMs to automate and innovate within the algorithm design process could lead to more efficient, effective, and creative algorithms for complex problems across various domains. We hope this paper can contribute to a deeper understanding of this potential and set the stage for future innovations and collaborations in this promising avenue of research.